### v8

# Structured Packing in LLM Training Improves Long Context Utilization

Abstract
Recent advancements in long-context large language models have attracted significant attention, yet their practical applications often suffer from suboptimal context utilization. This study investigates structuring training data to enhance semantic interdependence, demonstrating that this approach effectively improves context utilization. To this end, we introduce the Structured Packing for Long Context (SPLiCe) method, which utilizes retrieval to collate mutually relevant documents into long and coherent training examples. We validate SPLiCe empirically across models of varying sizes — $3$B, $7$B, and $13$B — achieving improved performance in long-context tasks, such as Qasper and HotpotQA. Remarkably, even brief fine-tuning with SPLiCe is sufficient to realize these benefits. Additionally, SPLiCe effectively mitigates the ‘lost-in-middle’ phenomenon often observed in large models. Our comprehensive analysis of SPLiCe explores its design choices and reveals intriguing transfer effects; for instance, training on programming code enhances performance on natural language tasks. 11footnotetext: University of Warsaw, 2Ideas NCBR, 3University of Edinburgh, 4Google DeepMind, 5Institute of Mathematics, Polish Academy of Sciences, 6deepsense.ai
Accepted to AAAI’25
1 Introduction
Large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023b; Bai et al., 2023) have transformed the field of AI. Recently, the field has observed the rise of long context Language Models (LCLMs) that promise to unveil novel and powerful capabilities (Anthropic, 2023; OpenAI, 2023a; Gemini Team, 2024). However, their ability to process long context is not always as effective as one hopes for. Indeed, several studies have highlighted an important limitation: when processing prompts composed of multiple documents, LCLMs frequently encounter difficulties in accurately extracting relevant information (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023a). Additionally, they typically find it challenging to utilize information from the middle of their inputs (Liu et al., 2023), even on simple synthetic retrieval tasks (Li et al., 2023a). Understanding these issues is vital for advancements in LCLM technologies and calls for systematic research.
In this work, we take a step towards better context utilization in LCLMs. We focus on training data, keeping other components, such as the architecture and training objectives, unchanged. The broad question is how to organize training data to enhance long context capabilities? Such perspective has received some attention recently (Levine et al., 2022; Chan et al., 2022; Shi et al., 2023b), but the problem is far from being solved. The central finding of this work is that structuring training data to increase semantic interdependence is an effective strategy towards better long context utilization. We achieve this by introducing and evaluating Structured Packing for Long Context (SPLiCe), a method for creating training examples by using retrieval (e.g., BM25, Contriver or repository structure) to collate mutually relevant documents into a single training context.
Figure 1: Training examples generated by the standard sampling procedure vs proposed by SPLiCe. Similar colors indicate related documents, which could be found using a retrieval method (e.g., BM25 or contriver) or metadata (e.g., repository structure for code).
We empirically validate SPLiCe showing that fine-tuning of OpenLLaMA $3$Bv2, $7$Bv2 (Geng and Liu, 2023) and CodeLlama $13$B (Rozière et al., 2023) with mere $2$B–$6$B tokens already brings substantial improvements in handling long context information in downstream tasks that require retrieval and in-context learning. These tasks include Qasper (Dasigi et al., 2021) from SCROLLS (Shaham et al., 2022), lost-in-the-middle benchmark (Liu et al., 2023), HotPotQA (Yang et al., 2018), TREC (Li and Roth, 2002; Hovy et al., 2001), and DBpedia (Lehmann et al., 2015). We perform a comprehensive study of the design choices and properties of SPLiCe, showing, in particular, that the acquired long context capabilities transfer between modalities, such as code and text. SPLiCe also helps to retain and in some cases even improve performance on short context benchmarks like GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021) and TREC (Li and Roth, 2002; Hovy et al., 2001).
Our contributions can be summarized as follows:
  • We comprehensively show that structuring training data is a viable way of improving the long context utilization of LLMs. To this end, we introduce SPLiCe, a method for creating training examples by using retrieval to collate mutually relevant documents into a single training context.
  • We fine-tune OpenLLaMA $3$Bv2, OpenLLaMA $7$Bv2 (Geng and Liu, 2023) and CodeLlama $13$B (Rozière et al., 2023) using SPLiCe, showing that it improves downstream performance on long context tasks.
  • We present a thorough analysis of the design choices behind SPLiCe, such as the number of retrieved documents and the order in which documents are merged into a training example.
2 Method
SPLiCe is a method for constructing training examples that enhance the LLMs’ long context utilization. This translates to improved performance on a range of tasks like in-context learning, question answering, in-context information retrieval, and long-context language modeling.
Algorithm 1 SPLiCe training example construction
Input:
   $D$: document corpus
   $k$: breadth hyper-parameter
   $L$: maximum length of returned training example
   RETRIEVE: retrieval method to use, e.g., BM25
   ORDER: ordering method, e.g., identity, or random shuffle
Output: training example consisting of concatenated documents
$d_{r}\sim D$Sample the root document}
$D=D\setminus\{d_{r}\}$
$C=[d_{r}]$
$Q=\texttt{empty queue }$
$Q.\texttt{PUSH}(d_{r})$
while $Q\neq\emptyset$ and $\texttt{len}(C)\leq L$ do
   $d=Q.\texttt{POP}()$
   $d_{1},\ldots,d_{k}=\texttt{RETRIEVE}(d,k)$
   {Retrieve top-$k$ most similar documents to $d$ using a selected method, e.g., BM25}
   for each $d_{i}$ in $d_{1},\ldots,d_{k}$ do
      if $d_{i}\in D$ then
         {RETRIEVE uses a precomputed index and may return documents that are already in $C$}
         $C=C.\texttt{APPEND}(d_{i})$Append $d_{i}$ to $C$}
         $Q.\texttt{PUSH}(d_{i})$
         $D=D\setminus\{d_{i}\}$
      end if
   end for
end while
return $\texttt{TRIM}(\texttt{ORDER}(C),L)$
Rationale and intuitions
Capturing long-range dependencies is believed to enhance language modeling and retrieval-augmentation (Borgeaud et al., 2022). However, it is an open question how to achieve such benefits in pre-training or fine-tuning. The primary difficulty comes from long-range dependencies being rare in training data (de Vries, 2023) and diminishing with distance. Thus, it is unlikely that a model will learn to utilize long context without more explicit guidance.
Recent studies indicate that structuring data, i.e., going beyond the i.i.d. paradigm, might be beneficial or even necessary to achieve good long-context performance. (Levine et al., 2022) develops a theory showing that the trained model establishes stronger dependencies between text segments in the same training example. Chan et al. (2022) indicates that specific distributional properties of the training data have a significant impact on the model’s in-context performance. Finally, concurrently to our work (Shi et al., 2023b) shows that pre-training on structured data improves the performance of LLMs.
SPLiCe follows these intuitions. Namely, it constructs training examples by collating mutually relevant documents to increase the dependency density, thus allowing the model to learn to utilize long context. We provide additional insights into the distributional properties of SPLiCe created data at the end of this section and in Appendix K.
Structured Packing for Long Context (SPLiCe)
SPLiCe starts by picking a random document from the dataset to create a root of a tree and continues in a breadth-first manner, each time appending top-$k$ similar documents from the corpus. The final sequence consists of the tree flattened according to a certain tree traversal strategy; see Algorithm 1.
The hyperparameter $k$ introduces flexibility, enabling interpolation between different modes of retrieval. Specifically, when $k=1$, SPLiCe simulates a long document by creating a path of related examples. For larger $k$ values, SPLiCe generates examples akin to those used in retrieval-augmented models, e.g., (Borgeaud et al., 2022). It is essential for the dataset to be well-curated to ensure the efficient functioning of SPLiCe. For instance, maintaining a low duplication rate encourages the model to leverage long context beyond merely copying. In this work, we employ the StarCoder (Li et al., 2023b) dataset, which has been deduplicated using the pipeline described in (Allal et al., 2023).
Many possible retrieval methods can be used with SPLiCe (Retrieve function in Algorithm 1). In our experiments, we test the following:
  • SPLiCe Repo : based on additional meta-information about the data, that is the repository structure of the code (Repo): we concatenate files using a depth-first search algorithm on the directory structure, that is files from the same directory are grouped together. A similar method has been pioneered by (Wu et al., 2022) and proposed in (Shi et al., 2023b) as an interesting future direction.
  • SPLiCe BM25 : based on BM25 (Robertson and Zaragoza, 2009; Bassani, 2023), a standard retrieval method that uses a bag-of-words approach to rank documents based on their similarity to a query.
  • SPLiCe Cont : based on Contriever-MSMARCO (Cont) (Izacard et al., 2022), a recently proposed retrieval method that uses a transformer to rank documents based on their similarity to a query.
Baselines One standard approach, commonly used in LLM training pipelines, is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context. This is known as example packing (Brown et al., 2020). Another popular approach restricts sampled documents to the same meta-class (for example, Wikipedia article, source code) (Groeneveld et al., 2024; Zhao et al., 2024). We compare SPLiCe against both.
SPLiCe computational efficiency Given the dataset sizes used in training LLMs, computational efficiency plays a non-trivial role. SPLiCe Repo is the fastest and easiest to implement but requires additional directory structure, i.e., it does not apply to general web data. SPLiCe BM25 uses a bag of words BM25 method that lacks deeper semantic encoding. SPLiCe Cont requires calculating embeddings for each document and retrieval based on the vector inner-product. The latter step can be done efficiently using fast approximate inner-product search, e.g., Faiss (Johnson et al., 2017).
Burstiness SPLiCe conjecturally falls into the framework presented in Chan et al. (2022), which shows that the distributional properties of the training data affect the in-context capabilities of transformer models. In particular, it indicates the importance of "burstiness", i.e., a flatter frequency distribution with a relatively higher mass on the rare, long-tail tokens appearing in a sequence. In Appendix K, we show that SPLiCe increases the burstiness of the training data (measured in terms of Zipf’s coefficient of token frequency) compared to the example packing baseline.
3 Experiments
In this section, we show that SPLiCe improves the long context performance of large-scale language models. To this end, we use $3$B, $7$B, and $13$B parameter models and evaluate on tasks that test in-context learning (Section 3.2.1), question answering (Section 3.2.2), and in-context information retrieval capabilities (Section 3.2.3). Additionally, we provide perplexity results in Appendix E. Next, we show that SPLiCe can improve the core model capabilities by testing the short context performance, see Section 3.2.5. Finally, in Section 3.3, we tune over 30 medium ($270$M parameters) models using different data mixtures and SPLiCe parameters, to analyze its design choices.
An important finding of our work is that presented improvements occur during a relatively short, compared to pre-training, fine-tuning. To be more precise, $3$B models were tuned on $5.4$B tokens, whereas $7$B and $13$B models were tuned on $2$B tokens. The $3$B and $7$B start from OpenLLaMAv2 (Geng and Liu, 2023) that was pre-trained on $1$T tokens, the $13$B parameter models start from CodeLlama (Rozière et al., 2023) that was pre-trained on $2.5$T tokens.
3.1 Experimental setup
For $3$B experiments, we continue training on a 50/50 mixture of RedPajama, prepared in the standard way, and C prepared using SPLiCe BM25 . For $7$B and $13$B ones, we continue the training on a 50/25/25 mixture of RedPajama (50) prepared in the standard way, StackExchange (25) and C (25) prepared using SPLiCe BM25 . StackExchange is part of RedPajama TogetherComputer (2023), and C data come from StarCoder (Li et al., 2023b). Including the standard RedPajama aims to prevent the model from overfitting to artificially created long documents. We refer to Appendix A for a detailed description of the model architectures.
We train with $32$K context length. We use a batch size of $256$K tokens per step and the learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng and Liu, 2023). For $13$B model we use batch of $512$K tokens per step.
Regarding the context extensions method, we use FoT (Tworkowski et al., 2023), and the CodeLlama (CL) (Rozière et al., 2023) approach. More explicitly, we test eight models:
[…]
where Baseline denotes the standard (example packing) data preparation method where context is created by sampling random documents from the corpus and separating them with BOS/EOS tokens. We provide results regarding the within-domain sampling baseline in Appendix C.2. If not stated otherwise, in SPLiCe we use $k=1$ and the identity permutation as Order in the Algorithm 1. Hyperparameter details can be found in Appendix A.
3.2 Experimental results
3.2.1 In-Context learning
To evaluate the improvements of in-context learning abilities, we use TREC (Li and Roth, 2002; Hovy et al., 2001) and DBpedia (Lehmann et al., 2015), which are text classification tasks. For TREC, to test the long context utilization, we vary the number of in-context examples in $\{780,1560\}$ (which correspond roughly to $\{16,32\}$K context length). We sample these examples multiple times and measure classification accuracy on the test set. In Table 1, we observe that SPLiCe improves the average performance across different context lengths and model sizes. We follow a very similar protocol for DBpedia ($\{16,32\}$K context length corresponding to $\{190,380\}$ in-context examples), confirming that SPLiCe improves the context utilization. In Appendix I, we study the distribution of improvements with respect to the choice of the in-context examples, showing the stochastic domination of SPLiCe over the baseline.
Table 1: We test the average classification performance on TREC (Li and Roth, 2002; Hovy et al., 2001). We average across $50$ sets of in-context examples for $3$B models, $10$ sets for $7$B models, and $5$ sets for $13$B models. We use $\Delta$ [confidence interval] to denote the mean improvement and its $95\%$ bootstrap confidence intervals (see Appendix I). For DBpedia, we average results across $40$ sets of in-context examples for $3$B and $7$B models and $5$ for $13$B ones. Due to the size of the DBpedia evaluation dataset, for each set of in-context examples, we subsample a subset of $500$ elements of the evaluation set.
Task TREC DBpedia
Model Context Baseline SPLiCe $\Delta$ [conf interv] Baseline SPLiCe $\Delta$ [conf interv]
[…]
$16$K 68.9 76.9 8.0 [6.9,9.3]{6.9,9.3]}6.9 , 9.3 ] 79.1 82.0 2.9 [2.5,3.4]{2.5,3.4]}2.5 , 3.4 ]
[…]
$16$K 74.0 79.0 5.0 [3.4,6.0]{3.4,6.0]}3.4 , 6.0 ] 83.6 85.6 2.0 [1.5,2.5]{1.5,2.5]}1.5 , 2.5 ]
[…]
$16$K 81.4 82.5 1.1 [0.2,1.6]{0.2,1.6]}0.2 , 1.6 ] 96.2 96.4 0.2 [0.0,0.3]{0.0,0.3]}0.0 , 0.3 ]
[…]
$16$K 88.2 91.2 3.0 [1.9,3.5]{1.9,3.5]}1.9 , 3.5 ] 95.8 96.8 1.0 [0.6,1.5]{0.6,1.5]}0.6 , 1.5 ]
Table 2: We measure question answering over long input using Qasper (Dasigi et al., 2021) (2-shot setting) and HotPotQA (Yang et al., 2018) (10-shot setting). For Qasper, we use the implementation from Language Model Evaluation Harness (Gao et al., 2021). For HotPotQA, we average results across $7$ sets of in-context examples for $3$B and $7$B models and $2$ for $13$B ones, with $\Delta$[confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence intervals. Note that in the $3$B model case, despite using SPLiCe for code data only, we still have improvements in non-code tasks.
Task QASPER HotPotQA
Model Context Baseline SPLiCe Context Baseline SPLiCe $\Delta$ [conf interv]
[…]
3.2.2 Question answering
We also show that SPLiCe models have improved question answering capabilities. To this end, we use Qasper (Dasigi et al., 2021) from SCROLLS (Shaham et al., 2022) and HotPotQA (Yang et al., 2018). Qasper contains questions regarding research papers provided as a part of the input context. In Table 2, we observe that SPLiCe improves the two-shot performance. Similarly, we observe improvements on HotPotQA, which tests the ability to aggregate knowledge across multiple documents. We stress that those results measure few-shot performance, as neither of the datasets was used during training.
3.2.3 Lost in the middle
The key retrieval task introduced in (Liu et al., 2023) has become a routine diagnostic tool to study context capabilities. Specifically, the model is presented with a list of key-value pairs, which are $128$-bit UUIDs, and is tasked to retrieve the value for a given key. Even though very simple, the task proved to be challenging for many open-source language models. Specifically, LMs tend to achieve better performance when the relevant information is either in the beginning or the end of the input context, while struggling in the middle. See Appendix D for the task details.
In Figure 2, we present the key retrieval performance of $7$B models trained with CodeLlama (Rozière et al., 2023) context extension method. We note that despite relatively short tuning, SPLiCe significantly helps on hard-to-retrieve positions (the position of the key in the range $[25,100]$). We present more details in Appendix D.
Figure 2: Key-value retrieval performance on a dictionary of $300$ key-value pairs ($\approx$ $24$K tokens). The $7$B CL model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions in the middle. The details about this task can be found in Appendix D. Each position was evaluated using $500$ examples.
3.2.4 Ordering of examples
We typically use the identity ordering as Order in Algorithm 1 to merge documents into a single context, as we found that in most cases, it performs best. In Appendix F, we found that random shuffling is slightly better in some cases. Specifically, this is the case of large $7$B CodeLlama models, see Table Table 18. We hypothesize that random ordering forces the model to make use of the full space of the RoPe positional encoding. Whereas the identity ordering, also used in (Shi et al., 2023b), skews the model toward paying more attention to fragments of text that are not too far away. This suggests an interesting research direction on the intersection of data preparation and positional embeddings.
3.2.5 Short context evaluation
An undesirable side effect of fine-tuning for extended context (e.g., with SPLiCe) could be performance degradation on shorter contexts. To test this, we evaluate the short context performance of our models using GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021) and TREC (Li and Roth, 2002; Hovy et al., 2001). We present the results in Table 3. We note that for smaller models, both SPLiCe and Baseline show some short context degradation, whereas the larger $13$B CodeLlama Model SPLiCe brings significant improvements in short context performance on GSM8K ($+1.7$), TREC ($+3.3$) and maintains the MMLU performance. We provide an additional evaluation on code in Appendix L.
Table 3: We evaluate our models on MMLU ($5$-shot), GSM8K ($8$-shot) and TREC ($90$-shot). We provide an additional comparison with their starting checkpoint. For $7$B case, we additionally compare with a model tuned with $2$k context length on the same data. For each task, we highlight best results up to $1$ point.
Model MMLU GSM8K TREC
STEM Humanities Social Other Mean
$3$B Starting Chkpt 24.4 25.6 26.3 27.8 25.9 4.3 58.4
[…]
$7$B Starting Chkpt 33.6 42.1 46.0 44.5 40.8 8.0 59.6
$7$B 2K Tuned 33.7 40.8 44.6 41.1 39.4 8.4 62.3
[…]
$13$B Starting Chkpt 36.6 48.2 50.6 45.4 44.3 21.4 78.7
[…]
3.3 Detailed study with medium models
In Table 5 and Table 5, we present a comprehensive examination of the impact of document packing on long-context performance using $270$M parameter models. Using the smaller size allows us to train over $30$ models and, thus, more accurately estimate the impact of various design choices.
Training protocol
Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer, 2023). Subsequently, we continue training using $1$B tokens with the context extended to $32$K on a mixture of the original RedPajama data (TogetherComputer, 2023) and long context data created using SPLiCe/Baseline. The amount of pre-training tokens is based on scaling laws from (Hoffmann et al., 2022) and constants calculated for GPT-like models in (Karpathy, 2022).
We employ the Focused Transformer (FoT) objective Tworkowski et al. (2023) for context extension (unless stated otherwise). This approach is motivated by practical factors, viz. training with short context length expedites the process, while context scaling can be achieved by finetuning on a relatively small amount of tokens, as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023). Loosely inspired by (Ouyang et al., 2022; Rozière et al., 2023), in the latter phase, long context data (i.e., prepared with SPLiCe) constitutes half of the mixture. We also check how SPLiCe works with other context extension methods in Appendix B.2, and 64K context length in Appendix J.
Evaluation
We measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets employing a context length of $32$K. The selection of these datasets is motivated by the fact that they can benefit from long-context information as demonstrated in (Chen et al., 2023; Li et al., 2023b). For example, functions are often re-utilized, and similar terms are employed across papers. We exclude documents with fewer than $32$K tokens and truncate those exceeding this length. In Appendix B.1 we evaluate our models using short context data confirming no performance degradation with respect to the base model. For information regarding the training and evaluation data, see Appendix G.
[…]
Altered Method arXiv Code Code &
Data Haskell Python CUDA All arXiv
C SPLiCe BM25 5.46 (.09) 3.20 (.17) 2.81 (.12) 2.22 (.11) 2.94 (.13) 3.10 (.13)
SPLiCe Cont 5.48 (.07) 3.22 (.15) 2.82 (.11) 2.23 (.10) 2.96 (.11) 3.11 (.12)
SPLiCe Repo 5.47 (.08) 3.22 (.15) 2.83 (.10) 2.24 (.09) 2.96 (.11) 3.12 (.11)
Baseline 5.55 3.37 2.93 2.33 3.07 3.23
C# SPLiCe BM25 5.52 (.13) 3.33 (.25) 2.90 (.17) 2.46 (.19) 3.11 (.20) 3.26 (.20)
SPLiCe Cont 5.53(.12) 3.35(.23) 2.91(.16) 2.48(.17) 3.12(.19) 3.27(.19)
SPLiCe Repo 5.53 (.12) 3.35 (.23) 2.91 (.16) 2.49 (.16) 3.12 (.19) 3.27 (.19)
Baseline 5.65 3.58 3.07 2.65 3.31 3.46
Python SPLiCe BM25 5.47 (.10) 3.25 (.21) 2.53 (.09) 2.41 (.15) 3.02 (.15) 3.17 (.15)
SPLiCe Cont 5.49 (.08) 3.28 (.18) 2.53 (.09) 2.43 (.13) 3.03 (.14) 3.19 (.13)
SPLiCe Repo 5.48 (.09) 3.27 (.19) 2.54 (.08) 2.44 (.12) 3.03 (.14) 3.18 (.14)
Baseline 5.57 3.46 2.62 2.56 3.17 3.32
Wikipedia SPLiCe BM25 5.64 (.09) 3.82 (.15) 3.26 (.11) 2.87 (.13) 3.55 (.13) 3.68 (.13)
SPLiCe Cont 5.65 (.08) 3.87 (.10) 3.30 (.07) 2.92 (.08) 3.59 (.09) 3.72 (.09)
Baseline 5.73 3.97 3.37 3.00 3.68 3.81
StackEX SPLiCe BM25 5.07 (.07) 3.88 (.06) 3.32 (.04) 2.89 (.05) 3.60 (.05) 3.69 (.05)
SPLiCe Cont 5.09 (.05) 3.91 (.03) 3.35 (.01) 2.93 (.01) 3.63 (.02) 3.73 (.01)
Baseline 5.14 3.94 3.36 2.94 3.65 3.74
Table 5: Perplexity fine-tune on a $/50$ 50 data mixture of RedPajama and C code. We report the mean and standard deviation. Interestingly, training on the code data with SPLiCe improves general long-context performance on arXiv.
Long Context Method arXiv Code Code &
Data Python All arXiv
C SPLiCe BM25 5.463 $\pm$ .002 2.810 $\pm$ .002 2.942 $\pm$ .005 3.100 $\pm$ .004
SPLiCe Cont 5.477 $\pm$ .005 2.824 $\pm$.001 2.957 $\pm$ .006 3.115 $\pm$ .006
SPLiCe Repo 5.474 $\pm$ .007 2.827 $\pm$ .006 2.958 $\pm$ .009 3.115 $\pm$ .009
Baseline 5.550 $\pm$ .002 2.931 $\pm$ .008 3.073 $\pm$ .006 3.228 $\pm$ .005
3.4 Summary of the results
Structuring data improves performance. We observe (see Tables 15) that all tested variants of SPLiCe outperform Baseline, often significantly, regardless of the training and eval dataset. This indicates that the structure of the data is important for long-context training.
The positive effects transfer between domains. We observe, see Tables 2 and 5, that the positive effects of SPLiCe transfer between domains; when training on code and evaluating on Qasper/arXiv and when training on web data and evaluating on code. This is a very interesting effect, which indicates that training induces generic long-context capabilities. Following that, we conjecture that better data mixtures and synthetic long-context data could bring further benefits.
SPLiCe improves the performance of various context extension methods SPLiCe only changes the input structure, and conjecturally, it should work with any context extension method. Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023). FoT extends the context only in a couple of attention layers and does not change the positional embeddings of the remaining layers. In Table 9 in Appendix B.2, we confirm that SPLiCe integrates seamlessly with methods that extend the context in all layers. Specifically, we test adjusting Rotary Positional Encodings according to either the recipe as in CodeLlama (Rozière et al., 2023), or YaRN (Peng et al., 2023), or left without changes.
Training is stable. We prepared three subsets of the C code dataset and ran the tested methods on each of them. In Table 5, we report the mean perplexity and its standard deviation. We observe that the differences between the subsets are minimal, which indicates training stability and confirms the statistical significance of our results.
4 Related work
There is an increasing number of works aiming to study the role of data in LLM training in detail. For instance, (Levine et al., 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training examples leads to better sentence embeddings and improves open-domain question-answering performance. Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include intrinsic tasks. They showed that this approach substantially boosts in-context learning. Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al., 2022). Work (Chan et al., 2022) identifies the training data’s distributional properties that affect transformer models’ in-context capabilities. Similarly, Han et al. (2023) constructs small-scale data using an iterative gradient approach and shows that such data improve in-context performance.
Our methodology diverges from these works in several key ways. Firstly, we focus on the document-level context during the training phase and long context performance, as opposed to sentence-level (Levine et al., 2022) or paragraph-level (Gu et al., 2023) granularity. We demonstrate the efficacy of our approach in large-scale language modeling, specifically with OpenLLaMA $3$B, $7$B, and CodeLlama $13$B. Secondly, we construct a tree structure of related documents through BM25/Contriever-MSMARCO retrieval and linearize this structure to form long-context examples. This allows for greater control over example coherence compared to relying solely on natural data structures like repository-level code. The gradient-based method presented in Han et al. (2023) can be broadly associated with retrieval used in our work; however, it differs in scale and granularity.
Concurrently to our research, (Shi et al., 2023b) showed a method for preparing the training data that closely resembles SPLiCe with default settings ($k=1$ and the identity as Order). The key differences are that (Shi et al., 2023b) focuses on training from scratch while we aim to achieve long context capabilities with short and cheap fine-tuning. Next, we use longer context lengths, up to $32$K tokens, whereas (Shi et al., 2023b) utilizes $8$K. Finally, we provide detailed analysis regarding the design choices, such as the benefits of data reordering in Section 3.2.4 and different values of $k$ in Appendix C.1.
5 Limitations and future work
We show that structuring the training data is a viable way of improving the model’s performance. The presented method, SPLiCe, can viewed as a general framework for organizing the documents into training examples. This opens multiple further research avenues.
Degree of relevance In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Cont/Repo). However, recent results of (Tworkowski et al., 2023) show that introducing unrelated data to the context may help the model learn better representations. We leave the study of how the choice of retriever (in particular, the ratio of related and unrelated documents) affects performance for future work.
In order to prepare the training data, SPLiCe uses each document exactly once (we mask out each used document for further retrieval). However, it is possible that allowing some high-quality documents to occur more than once may be beneficial.
Retrieval granularity Another avenue for future work is to study the granularity of the pieces from which the training examples are constructed. In this work, we focus on the document-level granularity. However, it is possible to construct training examples from smaller pieces.
Other context extension methods In most cases, our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023). Additionally, we tested three popular context extension methods on a medium scale (Naive, YaRN, and CodeLlama) and tuned a large $7$B model using the CodeLlama approach, preliminary confirming the applicability of SPLiCe. However, we leave more detailed studies with other context extension methods to future work.
Other data sources One of the approaches to training long-context language models is to use conversational data (Li et al., 2023a). This is complementary to our method. SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, or StackExchange questions and answers) to further increase the number of long-context examples. We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
Data curation Using highly correlated samples has the potential to result in training instability. However, we noted no performance degradation during our experiments. We leave the study of how SPLiCe integrates with different data types for the future.
6 Conclusions
In this work, we present SPLiCe, a method of constructing training examples for long-context language models. It utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner. We show that SPLiCe improves performance on downstream tasks and the language modeling abilities of LLMs. We further show that SPLiCe can be used to improve long-context utilization of large-scale models using only short fine-tuning. We believe that our work indicates multiple interesting research directions for improving the performance of long-context language models with structured data.
Acknowledgments and Disclosure of Funding
We gratefully acknowledge that our research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH, CI TASK) for providing computer facilities and support within computational grant numbers PLG/2023/016265. Piotr Miłoś was supported by National Science Center Poland under the grant agreement 2019/35/O/ST6/03464.
Appendix A Architecture
The architecture of our models is based on LLaMA [Touvron et al., 2023], and the architectural details can be found in Table 6. Briefly speaking, our architecture is similar to the one introduced in [Vaswani et al., 2017] with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform RMSNorm before the input of both the attention and feed-forward modules. Thirdly, we use the LLaMA FeedForward module. Additionally, we use Rotary Position Embedding [Su et al., 2021]. For context extension, we use Focused Transformer (FoT)[Tworkowski et al., 2023], CodeLlama [Rozière et al., 2023] (CL) and YaRN Peng et al. [2023]. Table 7 presents the details about both standard and long-context pretraining. We use AdamW as an optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.95$.
Table 6: Architecture details. Focused Transformer context extension is applied in continued pretraining for $32$K context and evaluation.
Parameter/Model Size 270M 3B 7B 13B
Vocab Size 32000 32000 32000 32016
Embedding Size 1024 3200 4096 5120
Num Attention Layers 12 26 32 40
Num Attention Heads 8 32 32 40
Head Size 128 100 128 128
MLP Hidden Size 4096 8640 11008 13824
FoT Context Extension Layers [6, 9] [6, 12, 18] [8, 16, 24] [12, 24, 36]
Each tuning experiment of 270M model was done using either a TPUv3-8 or TPUv4-8 machine and took around $10$ hours. Tuning a $3$B parameter model for $5.4$B tokens with FoT took $40$ hours on TPUv3-128. The $7$B and $13$B models were tuned on TPUv3-128.
Table 7: Training details. We pretrain a custom $270$M parameter model and take a pretrained $3$B/$7$B/ parameter OpenLLaMAv2 model [Geng, 2023] and $13$B CodeLlama [Rozière et al., 2023] model. Subscript denotes that parameter was specific for a context extension method with FoT referring to [Tworkowski et al., 2023] and no-FoT to other methods (Naive, YaRN [Peng et al., 2023], CodeLlama [Rozière et al., 2023]).
Stage Parameter/Model Size 270M 3B 7B 13B
Pretraining Context 2K 2K 2K 16K
Tokens 6.3B 1T 1T 2.5T
Long Context Tuning Context 32KFoT, 16Kno-FoT 32K 32K 32K
Batch Size 128FoT, 16no-FoT 128 128FoT, 32no-FoT 128
Start Learning Rate 5e-5 1.5e-5 1.5e-5 1.5e-5
End Learning Rate 5e-6 1.5e-6 1.5e-6 1.5e-6
Warmup Steps 250 1000 1000FoT, 200no-FoT 1000
Tokens 1B 5.4B 2B 2B
Appendix B Additional results for $270$M models
B.1 Short context evaluation
In Table 8, we asses the short context performance of $270$M models from Section 3.3 and compare them against their starting checkpoint.
Table 8: $2$K context perplexity evaluation of models from Section 3.3
Tune Dataset Method arXiv Code Code &
Haskell Python CUDA All arXiv
Starting Chkpt 10.80 13.40 6.00 5.00 7.17 7.40
Wikipedia SPLiCe BM25 10.64 13.28 6.00 5.00 7.16 7.38
SPLiCe Cont 10.60 13.24 5.99 4.99 7.14 7.36
Baseline 10.59 13.20 5.99 5.00 7.12 7.34
Stackexchange SPLiCe BM25 9.28 12.94 5.91 4.89 7.05 7.19
SPLiCe Cont 9.26 12.88 5.89 4.88 7.04 7.18
Baseline 9.24 13.00 5.91 4.90 7.07 7.21
C SPLiCe BM25 10.76 12.10 5.55 3.78 6.24 6.52
SPLiCe Cont 10.76 12.06 5.53 3.76 6.21 6.50
Baseline 10.73 11.96 5.49 3.68 6.16 6.44
SPLiCe Repo 10.78 12.10 5.56 3.79 6.24 6.52
C# SPLiCe BM25 10.86 12.74 5.73 4.67 6.72 6.98
SPLiCe Cont 10.89 12.72 5.72 4.65 6.70 6.96
Baseline 10.83 12.73 5.70 4.65 6.69 6.95
SPLiCe Repo 10.91 12.77 5.74 4.68 6.72 6.98
Python SPLiCe BM25 10.75 12.33 4.12 4.44 6.38 6.65
SPLiCe Cont 10.76 12.28 4.09 4.43 6.36 6.64
Baseline 10.72 12.19 3.98 4.40 6.33 6.61
SPLiCe Repo 10.78 12.33 4.12 4.45 6.38 6.66
B.2 Different context extension methods
We test SPLiCe with different context extension methods. In Table 9, we show that SPLiCe brings improvements also when context is extended in all layers using the naive approach, the CodeLlama [Rozière et al., 2023], and [Peng et al., 2023] RoPe [Su et al., 2021] adjustment method.
[…]
RoPe Method arXiv Code Code &
scale Haskell Python CUDA All arXiv
Naive SPLiCe BM25 6.25 (.08) 4.84 (.19) 3.55 (.11) 2.84 (.12) 3.72 (.13) 3.88 (.12)
SPLiCe Repo 6.25 (.08) 4.87 (.16) 3.56 (.10) 2.85 (.11) 3.74 (.11) 3.89 (.11)
Baseline 6.33 5.03 3.66 2.96 3.85 4.00
CL SPLiCe BM25 5.74 (.02) 4.28 (.09) 3.22 (.05) 2.53 (.05) 3.34 (.06) 3.49 (.06)
SPLiCe Repo 5.74 (.02) 4.28 (.09) 3.22 (.05) 2.54 (.04) 3.35 (.05) 3.50 (.05)
Baseline 5.76 4.37 3.27 2.58 3.40 3.55
YaRN SPLiCe BM25 5.77 (.02) 4.32 (.10) 3.24 (.05) 2.55 (.06) 3.37 (.07) 3.52 (.06)
SPLiCe Repo 5.77 (.02) 4.32 (.10) 3.24 (.05) 2.56 (.05) 3.38 (.06) 3.53 (.05)
Baseline 5.79 4.42 3.29 2.61 3.44 3.58
B.3 Detailed results
In this section, we extend results presented in Section 3.3. Details about the number of tokens used for evaluation can be found in Table 19.
Tables 10 and 11 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method. Table 10 contains detailed results from training on C# and Python. Table 11 contains results on C averaged across three different subsets of C (for details about the construction of those subsets, see Appendix G). Both tables show that SPLiCe outperforms the Baseline by a significant margin.
The main advantage of SPLiCe BM25 /SPLiCe Cont over the SPLiCe Repo approach is that it can also be used for non-structured data. Table 12 shows the detailed results of applying the SPLiCe on non-code data. Note that training on non-code data allows us to improve the model perplexity on the arXiv dataset in comparison to the model trained on code.
Tables 13, 14 consider models trained with YaRN [Peng et al., 2023], CodeLlama [Rozière et al., 2023] and Naive (no adjustment to RoPE) context extension methods. Table 15 shows that a simple artificial extension of example length via random concatenation of documents (C source files) does not help.
Table 10: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in the mentioned ways. For details about training, please refer to Section 3.3.
Altered Train Data: C#
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline SPLiCe Repo
ArXiv 5.52 5.53 5.65 5.53
C 2.39 2.40 2.50 2.40
C++ 2.60 2.61 2.74 2.62
CUDA 2.46 2.48 2.65 2.49
C# 1.82 1.82 1.90 1.82
Common Lisp 3.41 3.44 3.72 3.42
Dart 2.14 2.16 2.31 2.16
Emacs Lisp 8.41 8.46 8.85 8.41
Erlang 2.80 2.80 2.95 2.81
Fortran 3.68 3.71 4.05 3.72
Go 1.94 1.95 2.06 1.96
Groovy 3.01 3.03 3.25 3.04
Haskell 3.33 3.35 3.58 3.35
Java 2.09 2.10 2.22 2.10
Pascal 3.61 3.62 3.80 3.60
Python 2.90 2.91 3.07 2.91
Mean 3.26 3.27 3.46 3.27
Altered Train Data: Python
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline SPLiCe Repo
ArXiv 5.47 5.49 5.57 5.48
C 2.38 2.39 2.45 2.39
C++ 2.61 2.63 2.71 2.63
CUDA 2.41 2.43 2.56 2.44
C# 1.98 1.99 2.06 2.00
Common Lisp 3.23 3.28 3.47 3.27
Dart 2.15 2.17 2.27 2.17
Emacs Lisp 7.94 7.98 8.28 7.93
Erlang 2.70 2.71 2.82 2.71
Fortran 3.42 3.46 3.72 3.46
Go 1.95 1.96 2.03 1.96
Groovy 2.97 2.99 3.13 2.99
Haskell 3.25 3.28 3.46 3.27
Java 2.12 2.12 2.20 2.13
Pascal 3.57 3.58 3.73 3.58
Python 2.53 2.53 2.62 2.54
Mean 3.17 3.19 3.32 3.18
Table 11: To check the statistical significance of our results, we prepare three subsets of C (details in Appendix G and train the models on a 50/50 mixture of RedPajama data (organized in the standard way) and C data organized using one of the methods. Note that the standard deviation is much lower than the perplexity improvements from using SPLiCe.
Altered Train Data: C
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline SPLiCe Repo
ArXiv 5.463 $\pm$ 0.002 5.477 $\pm$ 0.005 5.550 $\pm$ 0.002 5.474 $\pm$ 0.007
C 2.126 $\pm$ 0.020 2.134 $\pm$ 0.020 2.173 $\pm$ 0.023 2.135 $\pm$ 0.020
C++ 2.396 $\pm$ 0.004 2.403 $\pm$ 0.002 2.467 $\pm$ 0.001 2.403 $\pm$ 0.006
CUDA 2.219 $\pm$ 0.002 2.235 $\pm$ 0.005 2.330 $\pm$ 0.009 2.239 $\pm$ 0.004
C# 1.939 $\pm$ 0.000 1.948 $\pm$ 0.001 2.016 $\pm$ 0.004 1.949 $\pm$ 0.002
Common Lisp 2.994 $\pm$ 0.072 3.042 $\pm$ 0.091 3.195 $\pm$ 0.078 3.043 $\pm$ 0.102
Dart 2.141 $\pm$ 0.002 2.155 $\pm$ 0.004 2.268 $\pm$ 0.009 2.156 $\pm$ 0.002
Emacs Lisp 7.857 $\pm$ 0.017 7.851 $\pm$ 0.020 8.098 $\pm$ 0.027 7.840 $\pm$ 0.019
Erlang 2.665 $\pm$ 0.002 2.680 $\pm$ 0.003 2.769 $\pm$ 0.007 2.676 $\pm$ 0.003
Fortran 3.306 $\pm$ 0.010 3.335 $\pm$ 0.010 3.554 $\pm$ 0.010 $3$$\pm$ 0.012
Go 1.910 $\pm$ 0.001 1.924 $\pm$ 0.007 1.999 $\pm$ 0.002 1.924 $\pm$ 0.006
Groovy 3.009 $\pm$ 0.001 3.026 $\pm$ 0.006 3.147 $\pm$ 0.013 3.025 $\pm$ 0.007
Haskell 3.198 $\pm$ 0.001 3.221 $\pm$ 0.001 3.371 $\pm$ 0.008 3.220 $\pm$ 0.008
Java 2.075 $\pm$ 0.002 2.086 $\pm$ 0.001 2.161 $\pm$ 0.006 2.085 $\pm$ 0.005
Pascal 3.492 $\pm$ 0.026 3.496 $\pm$ 0.019 3.622 $\pm$ 0.021 3.513 $\pm$ 0.014
Python 2.810 $\pm$ 0.002 2.824 $\pm$ 0.001 2.931 $\pm$ 0.008 2.827 $\pm$ 0.006
Mean 3.100 $\pm$ 0.004 3.115 $\pm$ 0.006 3.228 $\pm$ 0.005 3.115 $\pm$ 0.009
Table 12: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and other data organized using one of the methods. For details about training, please refer to Section 3.3. Note that the model trained with SPLiCe on StackExchange outperforms the one trained on code on arXiv evaluation, showing the benefits of SPLiCe’s applicability to non-code data.
Altered Train Data: StackExchange
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline
ArXiv 5.07 5.09 5.14
C 2.68 2.69 2.70
C++ 3.02 3.04 3.06
CUDA 2.89 2.93 2.94
C# 2.27 2.28 2.29
Common Lisp 4.02 4.06 4.08
Dart 2.58 2.60 2.61
Emacs Lisp 9.55 9.67 9.69
Erlang 3.13 3.16 3.18
Fortran 4.28 4.34 4.38
Go 2.24 2.25 2.27
Groovy 3.62 3.66 3.68
Haskell 3.88 3.91 3.94
Java 2.43 2.45 2.45
Pascal 4.08 4.11 4.14
Python 3.32 3.35 3.36
Mean 3.69 3.73 3.74
Altered Train Data: Wikipedia
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline
ArXiv 5.64 5.65 5.73
C 2.65 2.67 2.71
C++ 2.98 3.01 3.07
CUDA 2.87 2.92 3.00
C# 2.22 2.24 2.29
Common Lisp 3.87 3.96 4.08
Dart 2.51 2.55 2.61
Emacs Lisp 9.38 9.45 9.63
Erlang 3.13 3.16 3.23
Fortran 4.23 4.32 4.49
Go 2.18 2.21 2.26
Groovy 3.49 3.55 3.67
Haskell 3.82 3.87 3.97
Java 2.39 2.41 2.46
Pascal 4.32 4.23 4.40
Python 3.26 3.30 3.37
Mean 3.68 3.72 3.81
Table 13: Perplexity results comparing different ways of organizing the same data for non-FoT models. All runs started from the same $270M$ model with $2048$ context and were trained for $16$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C# code is organized in one of three ways.
Altered Train Data: C#
Context $16$K: CodeLlama
Eval/Method SPLiCe BM25 Baseline SPLiCe Repo
ArXiv 5.74 5.76 5.74
C 2.66 2.70 2.66
C++ 2.79 2.83 2.79
CUDA 2.53 2.58 2.54
C# 1.91 1.93 1.91
Common Lisp 3.78 3.85 3.79
Dart 2.28 2.33 2.28
Emacs Lisp 8.29 8.41 8.30
Erlang 3.57 3.64 3.58
Fortran 3.93 4.01 3.95
Go 1.99 2.03 2.00
Groovy 2.95 3.01 2.96
Haskell 4.28 4.37 4.28
Java 2.31 2.35 2.31
Pascal 3.67 3.72 3.67
Python 3.22 3.27 3.22
Mean 3.49 3.55 3.50
Context $16$K: YaRN
Eval/Method SPLiCe BM25 Baseline SPLiCe Repo
ArXiv 5.77 5.79 5.77
C 2.68 2.72 2.68
C++ 2.81 2.85 2.81
CUDA 2.55 2.61 2.56
C# 1.92 1.94 1.92
Common Lisp 3.83 3.92 3.84
Dart 2.30 2.36 2.30
Emacs Lisp 8.37 8.52 8.38
Erlang 3.60 3.67 3.61
Fortran 3.97 4.06 3.99
Go 2.00 2.04 2.01
Groovy 2.98 3.03 2.98
Haskell 4.32 4.42 4.32
Java 2.33 2.37 2.33
Pascal 3.69 3.75 3.69
Python 3.24 3.29 3.24
Mean 3.52 3.58 3.53
Table 14: Perplexity results comparing different ways of organizing the same data for non-FoT models. All runs started from the same $270M$ model with $2048$ context and were trained for $16$K context on a 50/50 mixture of RedPajama with a Naive context extension method.
Altered Train Data: C#
Context $16$K: Naive
Eval/Method SPLiCe BM25 Baseline SPLiCe Repo
ArXiv 6.25 6.33 6.25
C 2.88 2.96 2.89
C++ 3.04 3.13 3.05
CUDA 2.84 2.96 2.85
C# 2.04 2.08 2.04
Common Lisp 4.40 4.56 4.39
Dart 2.50 2.60 2.51
Emacs Lisp 9.25 9.46 9.25
Erlang 3.98 4.10 4.00
Fortran 4.56 4.79 4.59
Go 2.14 2.21 2.16
Groovy 3.27 3.39 3.28
Haskell 4.84 5.03 4.87
Java 2.52 2.60 2.53
Pascal 4.05 4.20 4.10
Python 3.55 3.66 3.56
Mean 3.88 4.00 3.89
Table 15: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C code is organized in one of three ways.
Altered Train Data: C
Eval/Method SPLiCe BM25 Baseline DomRnd
ArXiv 5.46 5.55 5.55
C 2.13 2.17 2.18
C++ 2.40 2.47 2.47
CUDA 2.22 2.33 2.33
C# 1.94 2.02 2.02
Common Lisp 2.99 3.20 3.18
Dart 2.14 2.27 2.27
Emacs Lisp 7.86 8.10 8.09
Erlang 2.67 2.77 2.77
Fortran 3.31 3.55 3.56
Go 1.91 2.00 2.00
Groovy 3.01 3.15 3.15
Haskell 3.20 3.37 3.37
Java 2.07 2.16 2.16
Pascal 3.49 3.62 3.64
Python 2.81 2.93 2.93
Mean 3.10 3.23 3.23
Appendix C Ablations
C.1 SPLiCe parameters
There are two important design choices related to SPLiCe. First, how many related documents are retrieved in each step (the parameter $k$ in Algorithm 1). Second, how the documents are ordered. Table 16 indicates that $k=1$ is the best choice, though the differences are rather small. We found that changing the order of documents in training examples hardly matters for $270$M models. We use ’standard’, as ordered by Algorithm 1, the reversed order, and random shuffling.
Table 16: Ablation of SPLiCe hyper-parameters. For each ablation, we have trained the same $270M$ parameter model using different data organization methods. Top-$k$ corresponds to the number of descendants chosen in the $\texttt{RETRIEVE}(d,k)$ step of the Algorithm 1. Reverse and shuffle correspond to the final order of examples $C$ returned by the Algorithm 1 (reverse – the order of documents in $C$ is reversed, shuffle – examples are shuffled.
Code
Method Top-$k$ Order C++ Haskell Python CUDA All
SPLiCe BM25 top 1 standard 2.38 3.20 2.82 2.23 2.93
reverse 2.38 3.21 2.82 2.23 2.93
top 2 standard 2.39 3.23 2.84 2.24 2.95
reverse 2.39 3.24 2.84 2.24 2.95
top 3 standard 2.39 3.24 2.84 2.25 2.97
reverse 2.39 3.25 2.84 2.25 2.96
shuffle 2.40 3.24 2.84 2.25 2.96
We also evaluated the influence of BOS and EOS tokens on the performance of trained models. As both Repo and SPLiCe methods concatenate documents to create training examples, they effectively slightly decrease the number of separating tokens compared to the Baseline. However, in Appendix C.2 we included experiments showing that this has no impact on model performance.
C.2 Importance of separating tokens and in-domain sampling
We also evaluate the influence of BOS and EOS tokens on the performance. To be more precise, in all setups, training examples are separated by BOS and EOS tokens. As SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens. To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens and concatenation of documents from the same domain, we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the C dataset (sampling without replacement). The results are shown in Table 17. We note that the difference between the Baseline and the random concatenation approach is small, and the random concatenation approach does not result in significant perplexity gains. When concatenating the documents, we concatenate them within the domain (C language in this case).
Table 17: Perplexity evaluation of two methods of organizing the data. Baseline – example packing, document equals training example. DomRnd – concatenate documents within the domain into examples of length bounded by $120$K characters. Training examples are then fed into the model and separated by BOS and EOS tokens. The difference is negligible, which suggests that the extension of example length in DomRnd does not help the model to utilize extended context. Experiments were performed using the $270$M parameter model. We performed three runs on different subsets of C to provide mean and standard deviation. The training data is a 50/50 mixture of RedPajama, organized in a standard way, and C data, organized using a method of choice.
Altered Method arXiv Code Code &
Data Python All arXiv
C DomRnd 5.554 $\pm$ 0.004 2.931 $\pm$ 0.003 3.076 $\pm$ 0.005 3.231 $\pm$ 0.005
Baseline 5.550 $\pm$ 0.002 2.931 $\pm$ 0.008 3.073 $\pm$ 0.006 3.228 $\pm$ 0.005
Appendix D Key-Value retrieval task
Figure 2 shows how training on SPLiCe organized data improves the performance on the key-value retrieval task proposed in [Liu et al., 2023]. This is a zero-shot task in which a model is prompted with a JSON dictionary and asked to retrieve a value corresponding to a specified key. The structure of the input is showcased below.
Extract the value corresponding to the specified key in the JSON object below. JSON data: {"4ef217b7-6bc0-48c6-af35-2765f1e730f3": "068192b7-16b1-40e0-8495-61c63f979d50",  "cd6b8bdc-bc6c-4490-acb4-bc187a2dccba": "7364a26e-289f-4968-93d3-b273e882bdee",  "7d057372-4ab8-4811-8110-658c3f19fff4": "3ad075c5-b567-4201-85a7-cb31a0c91540",  "c62e192d-45e6-4646-bb88-1529c73256c9": "f0411644-1f6d-42a6-8af8-f06da66efc77",  "06134e93-e158-490e-a66c-8e3b98e12735": "50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",  "3286f978-4270-4b54-8bfa-540d7e0772e6": "075cc716-1836-4f90-9be3-53e3d4ec6585",  "4701aa05-c523-4b89-9700-64ab9c37c537": "49d86354-74c4-4256-9b3a-35e6e2b80d00",  "c8895805-e574-4f13-9fe5-89da1d8c4748": "cc91af7f-8509-4bdc-bad7-2646af68e6d2"}  "4701aa05-c523-4b89-9700-64ab9c37c537":
We noted that FoT-trained models struggle with this task. This is probably due to the fact that they extend context only in a couple of layers, and the key-value retrieval task requires looking up and extracting a long sequence of letters and digits. Because of that, we evaluate FoT models with shorter dictionaries consisting of $75$ key-value pairs (around $6$K tokens) and show the results in Figures 3, and 3. For comparison, we also evaluate the $7$B CL model with this context length and show the results in Figure 3.
Figure 3: Performance on a smaller version of key-value retrieval task from [Liu et al., 2023]. We note that FoT models (a), (b) generally struggle to retrieve tokens that are only visible to a subset of layers with extended context. For comparison, we show the results with a model that has extended context in all layers (c) using CodeLlama [Rozière et al., 2023] method of context extension. Each position was evaluated using $500$ examples.
Appendix E Perplexity improvements
In Figure 4 we present perplexity improvements of $3$B FoT SPLiCe BM25 over Baseline. Figure 5 shows the evolution of SPLiCe model perplexity during the training. We follow [Anthropic, 2023] and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets. We average perplexity across arXiv, CUDA, Haskell, and CommonCrawl datasets.
Figure 4: Perplexity improvement with SPLiCe against the Baseline of the final models (after $21$k training steps). We bucket tokens by their positions in the document and calculate the average. Each dot is the difference of the averages of the SPLiCe and Baseline models. We observe that SPLiCe has smaller perplexity, and the improvements tend to be larger for tokens further in the document.
Figure 5: Evolution of the perplexity with SPLiCe, as the model is trained on more tokens. See 4 for the difference with the baseline. As expected, SPLiCe significantly improves perplexity for tokens whose positions are very distant in the sequence. Perplexity for more distant tokens improves more significantly compared to tokens in the beginning, early in the training.
Appendix F Shuffling
Table 18: Average classification performance on TREC [Li and Roth, 2002, Hovy et al., 2001]. We compare $7$B CL model trained on SPLiCe prepared data with different approaches to ordering the examples (different function Order in 1). SPLiCe-shuf denotes the model trained on data that shuffled the documents randomly in context, for SPLiCe-no-shuf Order was the identity function. We use $\pm$ to denote the standard deviation. We decided to stick with the model trained with random shuffling as it has slightly better long-context performance and lower standard deviation.
TREC
Model
Context length
(# of examples)
SPLiCe-no-shuf SPLiCe-shuf
$7$B CL $32$K (1560) 76.0 $\pm$2.5 76.6 $\pm$1.9
$8$K (380) 77.1 $\pm$2.4 76.5 $\pm$1.8
Appendix G Data preparation
G.1 Evaluation data
We have taken a random subset of arXiv from Proof-pile. For StarCoder data, we have downloaded up to $64$GB of each of the mentioned language subsets and performed a random 85/15 split for languages that we train on.
When evaluating the perplexity of the model, we skip documents that are shorter than the model context and truncate documents that are longer than that. Table 19 shows the number of tokens over which the perplexity was calculated.
Table 19: Number of evaluation tokens in each of the considered datasets. For each context length $c$, we consider only documents that have not less than $c$ tokens and extract the $c$ tokens prefix.
Eval/Method 16K 32K 64K
ArXiv 16M 16M 16M
C 16M 16M 16M
C++ 16M 16M 16M
CUDA 16M 14M 8M
C# 16M 16M 16M
Common Lisp 16M 16M 16M
Dart 16M 16M 16M
Emacs Lisp 16M 15M 9M
Erlang 16M 16M 12M
Fortran 16M 16M 16M
Go 16M 16M 16M
Groovy 11M 4M 2M
Haskell 16M 16M 15M
Java 16M 16M 16M
Pascal 16M 16M 16M
Python 16M 16M 16M
G.2 Train data
The StackExchange data was taken from the Proof-pile. To prepare the code train data, we take the StarCoder train splits mentioned in Section G.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs. We also split repos larger than $25$MB and filter out files that are longer than $30$k characters. The reason behind repo splitting is to avoid the situation where one repository occupies a significant portion of the data pack. We have noticed repos containing as many as $40$K files and files as long as $11$M characters. The character filtering is consistent with our method as we aim to improve the performance in a scenario that lacks high-quality long-context data. For C# and Python, only one pack is used to organize the data. For C, we have performed a run on three packs and provided results and standard deviation in Table 5. For large models, we run the methods on several packs and concatenate the results into a single dataset. For natural language datasets, we extract a random subset of documents.
Appendix H Faiss parameters
Our experiments with SPLiCe Cont utilize Faiss [Johnson et al., 2019] for fast approximate inner-product search. To be more precise, we use the "IVF8192,Flat" index that we train on 262144 examples coming from the dataset.
Appendix I Detailed accuracy improvements
The performance of in-context learning depends much on the choice of the in-context examples. To study this in more detail we study the following random variable
$\Delta(c)=\text{ACC}_{\textsc{{SPLiCe}}}(c)-\text{ACC}_{\textsc{Baseline}}(c),$
where $\text{ACC}_{\textsc{{SPLiCe}}}(c),\text{ACC}_{\textsc{Baseline}}(c)$ are the accuracies of the model trained with SPLiCe and Baseline respectively on the random choice of in-context examples $c$. Below we report the histograms of $\delta(c)$. In Table 1 and Table 2 we report mean $\Delta$ and $95\%$ confidence intervals as $\Delta$ [confidence interval].
Figure 6 shows additional details about accuracy improvements on TREC when considering different numbers of in-context examples.
(a) Examples: $190$, Context $4$K
(b) Examples: $380$, Context $8$K
(c) Examples: $780$, Context $16$K
(d) Examples: $1560$, Context $32$K
Figure 6: Histograms of accuracy improvement of SPLiCe BM25 over Baseline on TREC question classification task. The results are obtained by comparing the accuracy on the test set of TREC of the $3$B FoT model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples. Each set of in-context examples consists of elements randomly sampled (without replacement) from the training subset of TREC. Note that the model trained with SPLiCe is almost always better than the Baseline.
(a) Examples: $190$, Context $16$K, Model: $3$B FoT
(b) Examples: $380$, Context $32$K, Model: $3$B FoT
(c) Examples: $190$, Context $16$K, Model: $7$B FoT
(d) Examples: $380$, Context $32$K, Model: $7$B FoT
(e) Examples: $190$, Context $16$K, Model: $7$B CL
(f) Examples: $380$, Context $32$K, Model: $7$B CL
Figure 7: Histograms of accuracy improvement of SPLiCe BM25 over Baseline on DBPedia. We sample $40$ sets of in-context examples and, for each set of in-context examples, evaluate on a random $500$ element subset of the DBPedia test set.
Appendix J Longer context
In Table 20, we present the perplexity results for $270$M parameter models trained as described in Section 3.3 but with 2x larger context extension – context extended from $2$K to $64$K, instead of from $2$K to $32$K.
[…]
Altered Method arXiv Code Code &
Data Haskell Python CUDA All arXiv
C# SPLiCe BM25 4.86 (.15) 2.60 (.19) 2.66 (.16) 2.32 (.19) 2.75 (.19) 2.88 (.19)
SPLiCe Cont 4.88 (.13) 2.62 (.17) 2.67 (.15) 2.34 (.17) 2.77 (.17) 2.90 (.17)
SPLiCe Repo 4.88 (.13) 2.62 (.17) 2.68 (.14) 2.35 (.16) 2.77 (.17) 2.90 (.17)
Baseline 5.01 2.79 2.82 2.51 2.94 3.07
Appendix K Data distribution property
Chan et al. [2022], Han et al. [2023] shows that the "burstiness" property of training data benefits a model’s generalisation. Specifically, "burstiness" presents a flatter frequency distribution, with a relatively higher mass on the rare, long-tail tokens appearing in a sequence, which results in a lower Zipf’s coefficient of token frequency. We also observe a burstiness property of data in SPLiCe, as shown in Table 21.
Table 21: Zipf’s coefficient of token frequency on Baseline and SPLiCe. A lower Zipf’s coefficient represents a more significant burstiness property.
Training Data Context Length Method Zipf’s Coefficient
C 2K SPLiCe BM25 ${1.667}_{(0.139)}$
Baseline ${1.717}_{(0.152)}$
StackEx 2K SPLiCe BM25 ${1.836}_{(0.088)}$
Baseline ${1.865}_{(0.074)}$
C 32K SPLiCe BM25 ${1.512}_{(0.055)}$
Baseline ${1.593}_{(0.025)}$
StackEx 32K SPLiCe BM25 ${1.643}_{(0.026)}$
Baseline ${1.664}_{(0.013)}$
Appendix L HumanEval
Table 22: We additionally evaluate our $7$B and $13$B parameter models on HumanEval [Chen et al., 2021]. For $7$B models, all methods experience a decrease in performance, but SPLiCe is better at maintaining short context performance. For $13$B, both SPLiCe and Baseline increase the performance. We note that in our evaluation we do not use special tokens designed for CodeLlama $13$B code completion. Instead, evaluate all models in the same way. Because of the different evaluation pipeline our results for Starting Checkpoint are not the same as the ones presented in [Rozière et al., 2023].
Model Method Pass@1
$7$B Starting Checkpoint 12.9
[…]
Baseline 9.1
[…]
Baseline 8.5
$13$B Starting Checkpoint 23.8
[…]
Baseline 26.8

### v9

@@
 Abstract
-Recent advancements in long-context large language models have attracted significant attention, yet their practical applications often suffer from suboptimal context utilization.
-This study investigates structuring training data to enhance semantic interdependence, demonstrating that this approach effectively improves context utilization.
-To this end, we introduce the Structured Packing for Long Context (SPLiCe) method, which utilizes retrieval to collate mutually relevant documents into long and coherent training examples.
-We validate SPLiCe empirically across models of varying sizes — $3$B, $7$B, and $13$B — achieving improved performance in long-context tasks, such as Qasper and HotpotQA.
-Remarkably, even brief fine-tuning with SPLiCe is sufficient to realize these benefits.
-Additionally, SPLiCe effectively mitigates the ‘lost-in-middle’ phenomenon often observed in large models.
-Our comprehensive analysis of SPLiCe explores its design choices and reveals intriguing transfer effects; for instance, training on programming code enhances performance on natural language tasks.
-11footnotetext: University of Warsaw, 2Ideas NCBR, 3University of Edinburgh, 4Google DeepMind, 5Institute of Mathematics, Polish Academy of Sciences, 6deepsense.ai
-Accepted to AAAI’25
+Recent advancements in long-context language modeling have attracted significant attention, yet their practical applications often suffer from suboptimal context utilization.
+To efficiently address this issue, we introduce the Structured Packing for Long Context, SPLiCe, a method that uses retrieval to collate mutually relevant documents into long training samples.
+We demonstrate that SPLiCe improves performance on long-context tasks, particularly by achieving perfect accuracy on the synthetic Needle in the Haystack benchmark, and effectively mitigating the ‘lost-in-the-middle’ phenomenon often observed in large language models.
+Notably, these long-context capabilities also extend to realistic downstream tasks, such as Qasper, across multiple model sizes—3B, 7B, and 13B—and are achieved with only brief fine-tuning on 2-6 billion tokens.
+We supplement these results with a detailed analysis of SPLiCe, examining the impact of hyperparameter choices, the different mixtures and proportions of SPLiCe-generated training data, and the choice of the retriever.
+We also study the transfer of long-context utilization skills between the modalities.
+An intriguing finding from our analysis is that training on a corpus of code can enhance performance on natural language tasks.
+Codehttps://github.com/ideas-ncbr/publications˙2024
+Extended versionhttps://arxiv.org/abs/2312.17296
 1 Introduction
-Large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023b; Bai et al., 2023) have transformed the field of AI.
-Recently, the field has observed the rise of long context Language Models (LCLMs) that promise to unveil novel and powerful capabilities (Anthropic, 2023; OpenAI, 2023a; Gemini Team, 2024).
-However, their ability to process long context is not always as effective as one hopes for.
-Indeed, several studies have highlighted an important limitation: when processing prompts composed of multiple documents, LCLMs frequently encounter difficulties in accurately extracting relevant information (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023a).
-Additionally, they typically find it challenging to utilize information from the middle of their inputs (Liu et al., 2023), even on simple synthetic retrieval tasks (Li et al., 2023a).
+Large language models (LLMs) (Brown et al.
+2020; Chowdhery et al.
+2022; Lewkowycz et al.
+2022; OpenAI 2023a; Bai et al.
+2023) have transformed the field of AI.
+Recently, the field has observed the rise of Long Context Language Models (LCLMs) that promise to unveil novel and powerful capabilities (Anthropic 2023; OpenAI 2023b; Gemini Team 2024).
+However, their ability to process long contexts is not always as effective as one hopes.
+Indeed, several studies have highlighted an important limitation: when processing prompts composed of multiple documents, LCLMs frequently encounter difficulties in accurately extracting relevant information (Tworkowski et al.
+2023; Liu et al.
+2023; Shi et al.
+2023; Kamradt 2023).
+Additionally, they typically find it challenging to utilize information from the middle of their inputs (Liu et al.
+2023), even on simple synthetic retrieval tasks (Li et al.
+2023a).
 Understanding these issues is vital for advancements in LCLM technologies and calls for systematic research.
@@
 We focus on training data, keeping other components, such as the architecture and training objectives, unchanged.
-The broad question is how to organize training data to enhance long context capabilities?
-Such perspective has received some attention recently (Levine et al., 2022; Chan et al., 2022; Shi et al., 2023b), but the problem is far from being solved.
+The broad question is: Given training data consisting of documents, how should these documents be organized into training samples to enhance long-context capabilities?
+While this perspective has received some attention recently (Levine et al.
+2022; Chan et al.
+2022; Shi et al.
+2024), the problem remains unsolved.
 The central finding of this work is that structuring training data to increase semantic interdependence is an effective strategy towards better long context utilization.
-We achieve this by introducing and evaluating Structured Packing for Long Context (SPLiCe), a method for creating training examples by using retrieval (e.g., BM25, Contriver or repository structure) to collate mutually relevant documents into a single training context.
-Figure 1: Training examples generated by the standard sampling procedure vs proposed by SPLiCe.
-Similar colors indicate related documents, which could be found using a retrieval method (e.g., BM25 or contriver) or metadata (e.g., repository structure for code).
-We empirically validate SPLiCe showing that fine-tuning of OpenLLaMA $3$Bv2, $7$Bv2 (Geng and Liu, 2023) and CodeLlama $13$B (Rozière et al., 2023) with mere $2$B–$6$B tokens already brings substantial improvements in handling long context information in downstream tasks that require retrieval and in-context learning.
-These tasks include Qasper (Dasigi et al., 2021) from SCROLLS (Shaham et al., 2022), lost-in-the-middle benchmark (Liu et al., 2023), HotPotQA (Yang et al., 2018), TREC (Li and Roth, 2002; Hovy et al., 2001), and DBpedia (Lehmann et al., 2015).
+We achieve this by introducing and evaluating Structured Packing for Long Context (SPLiCe), a method for creating training samples by using retrieval (e.g., BM25, Contriever) to collate mutually relevant documents into a single training context.
+We empirically validate SPLiCe showing that fine-tuning of OpenLLaMA $3$Bv2, $7$Bv2 (Geng and Liu 2023) and CodeLlama $13$B (Rozière et al.
+2023) with mere $2$B–$6$B tokens already brings improvements in handling long context information in downstream tasks that require retrieval and in-context learning.
+These tasks include Qasper (Dasigi et al.
+2021) from SCROLLS (Shaham et al.
+2022), HotPotQA (Yang et al.
+2018), Needle In A Haystack (Kamradt 2023), TREC (Li and Roth 2002; Hovy et al.
+2001), and DBpedia (Lehmann et al.
+2015).
+We also show that SPLiCe significantly alleviates the ’lost-in-the-middle’ phenomenon (Liu et al.
+2023) and outperforms standard example packing on the Needle In A Haystack task (Kamradt 2023) (see Figure 1).
 We perform a comprehensive study of the design choices and properties of SPLiCe, showing, in particular, that the acquired long context capabilities transfer between modalities, such as code and text.
-SPLiCe also helps to retain and in some cases even improve performance on short context benchmarks like GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021) and TREC (Li and Roth, 2002; Hovy et al., 2001).
+SPLiCe also helps to retain and in some cases even improve performance on short context benchmarks like GSM8K (Cobbe et al.
+2021), MMLU (Hendrycks et al.
+2021) and TREC (Li and Roth 2002; Hovy et al.
+2001).
+Figure 1: SPLiCe vs Example Packing (EP) (baseline) on Needle in a Haystack.
+A model fine-tuned with SPLiCe achieves perfect accuracy in retrieving fine-grained information over the whole context, while the baseline can only handle a small final segment (details in Appendix N).
 Our contributions can be summarized as follows:
-  • We comprehensively show that structuring training data is a viable way of improving the long context utilization of LLMs.
-To this end, we introduce SPLiCe, a method for creating training examples by using retrieval to collate mutually relevant documents into a single training context.
-  • We fine-tune OpenLLaMA $3$Bv2, OpenLLaMA $7$Bv2 (Geng and Liu, 2023) and CodeLlama $13$B (Rozière et al., 2023) using SPLiCe, showing that it improves downstream performance on long context tasks.
-  • We present a thorough analysis of the design choices behind SPLiCe, such as the number of retrieved documents and the order in which documents are merged into a training example.
+  • We comprehensively show that structuring training data is a viable way of improving the long context utilization.
+To this end, we introduce SPLiCe, a method for creating training samples by using retrieval to collate mutually relevant documents into a single sample.
+  • We fine-tune OpenLLaMA $3$Bv2, OpenLLaMA $7$Bv2 (Geng and Liu 2023) and CodeLlama $13$B (Rozière et al. 2023) using SPLiCe, showing that it improves long-contex downstream performance.
+  • We provide a comprehensive analysis of SPLiCe’s design choices, including retrieval parameters and document concatenation order, and evaluate its robustness and scalability with varying data sources and a parametrizable noisy retriever.
 2 Method
-SPLiCe is a method for constructing training examples that enhance the LLMs’ long context utilization.
-This translates to improved performance on a range of tasks like in-context learning, question answering, in-context information retrieval, and long-context language modeling.
-Algorithm 1 SPLiCe training example construction
-Input:
-   $D$: document corpus
-   $k$: breadth hyper-parameter
-   $L$: maximum length of returned training example
-   RETRIEVE: retrieval method to use, e.g., BM25
-   ORDER: ordering method, e.g., identity, or random shuffle
-Output: training example consisting of concatenated documents
+SPLiCe is a method for constructing training samples that improve the effectiveness of long-context fine-tuning.
+This leads to improved performance in tasks such as in-context learning, question answering, information retrieval, and long-context language modeling (see Section 3).
+Rationale and Intuitions
+Capturing long-range dependencies is believed to enhance language modeling and retrieval-augmentation (Borgeaud et al.
+2022).
+It is an open question how to achieve such benefits in pre-training or fine-tuning.
+The primary difficulty comes from long-range dependencies being rare in training data (de Vries 2023) and diminishing with distance.
+Thus, it is unlikely that a model will learn to utilize long context without more guidance.
+Recent studies indicate that structuring data, i.e., going beyond the i.i.d.
+paradigm, might be beneficial or even necessary to achieve good long-context performance.
+(Levine et al.
+2022) develops a theory showing that the trained model establishes stronger dependencies between text segments in the same training sample.
+Whereas concurrently to our work (Shi et al.
+2024) shows that pre-training on structured data improves performance (see also Section 4 for a more detailed comparison).
+SPLiCe follows these intuitions, and constructs training samples by concatenating mutually relevant documents to increase the dependency density, thus allowing the model to learn to utilize long context.
+Structured Packing for Long Context (SPLiCe)
+SPLiCe starts by picking a random document from the dataset to create a root of a tree and continues in a breadth-first manner, each time appending top-$k$ similar documents from the corpus.
+The final sequence is generated by flattening the tree according to a specific traversal strategy; see Algorithm 1.
+The hyperparameter $k$ introduces flexibility, enabling interpolation between different retrieval modes.
+Specifically, when $k=1$, SPLiCe simulates a long document by creating a path of related examples.
+For larger $k$ values, SPLiCe generates examples akin to those used in retrieval-augmented models, e.g.
+(Borgeaud et al.
+2022).
+SPLiCe Retrieval
+Many possible retrieval methods can be used with SPLiCe (Retrieve function in Algorithm 1). In our experiments, we test the following:
+  • SPLiCeRepo: based on additional meta-information about the data, that is the repository structure of the code (Repo): we concatenate files using a depth-first search algorithm on the directory structure, that is files from the same directory are grouped together.
+A similar method has been pioneered by (Wu et al.
+2022) and proposed in (Shi et al.
+2024) as an interesting future direction.
+  • SPLiCeBM25: based on BM25 (Robertson and Zaragoza 2009; Bassani 2023), a standard retrieval method that uses a bag-of-words approach to rank documents based on their similarity to a query.
+  • SPLiCeCont: based on Contriever-MSMARCO (Cont) (Izacard et al. 2022), a retrieval method that uses a transformer to rank documents based on their similarity.
+Figure 2: Training samples generated by Example Packing, Within-Domain Example packing, and SPLiCe.
+Similar colors and shapes indicate related documents, which could be found using a retrieval method (e.g., BM25 or Contriever) or metadata (e.g., git repository structure).
+Figure 3: Key-value retrieval performance on a dictionary of $300$ key-value pairs ($\approx$ $24$K tokens).
+The $7$B CL model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions in the middle than the Example Packing Baseline.
+The details about this task can be found in Appendix D.
+Each position averaged over $500$ examples.
+Algorithm 1 SPLiCe training sample construction
+  Input:    $D$: document corpus   $k$: breadth hyper-parameter   $L$: maximum length of returned training sample   RETRIEVE: retrieval method to use, e.g., BM25   ORDER: ordering method, e.g., identity, or shuffle
+  Output: training sample
   $d_{r}\sim D$Sample the root document}
@@
   end while
-return $\texttt{TRIM}(\texttt{ORDER}(C),L)$
-Rationale and intuitions
-Capturing long-range dependencies is believed to enhance language modeling and retrieval-augmentation (Borgeaud et al., 2022).
-However, it is an open question how to achieve such benefits in pre-training or fine-tuning.
-The primary difficulty comes from long-range dependencies being rare in training data (de Vries, 2023) and diminishing with distance.
-Thus, it is unlikely that a model will learn to utilize long context without more explicit guidance.
-Recent studies indicate that structuring data, i.e., going beyond the i.i.d.
-paradigm, might be beneficial or even necessary to achieve good long-context performance.
-(Levine et al., 2022) develops a theory showing that the trained model establishes stronger dependencies between text segments in the same training example.
-Chan et al.
-(2022) indicates that specific distributional properties of the training data have a significant impact on the model’s in-context performance.
-Finally, concurrently to our work (Shi et al., 2023b) shows that pre-training on structured data improves the performance of LLMs.
-SPLiCe follows these intuitions.
-Namely, it constructs training examples by collating mutually relevant documents to increase the dependency density, thus allowing the model to learn to utilize long context.
-We provide additional insights into the distributional properties of SPLiCe created data at the end of this section and in Appendix K.
-Structured Packing for Long Context (SPLiCe)
-SPLiCe starts by picking a random document from the dataset to create a root of a tree and continues in a breadth-first manner, each time appending top-$k$ similar documents from the corpus.
-The final sequence consists of the tree flattened according to a certain tree traversal strategy; see Algorithm 1.
-The hyperparameter $k$ introduces flexibility, enabling interpolation between different modes of retrieval.
-Specifically, when $k=1$, SPLiCe simulates a long document by creating a path of related examples.
-For larger $k$ values, SPLiCe generates examples akin to those used in retrieval-augmented models, e.g., (Borgeaud et al., 2022).
-It is essential for the dataset to be well-curated to ensure the efficient functioning of SPLiCe.
-For instance, maintaining a low duplication rate encourages the model to leverage long context beyond merely copying.
-In this work, we employ the StarCoder (Li et al., 2023b) dataset, which has been deduplicated using the pipeline described in (Allal et al., 2023).
-Many possible retrieval methods can be used with SPLiCe (Retrieve function in Algorithm 1). In our experiments, we test the following:
-  • SPLiCe Repo : based on additional meta-information about the data, that is the repository structure of the code (Repo): we concatenate files using a depth-first search algorithm on the directory structure, that is files from the same directory are grouped together.
-A similar method has been pioneered by (Wu et al., 2022) and proposed in (Shi et al., 2023b) as an interesting future direction.
-  • SPLiCe BM25 : based on BM25 (Robertson and Zaragoza, 2009; Bassani, 2023), a standard retrieval method that uses a bag-of-words approach to rank documents based on their similarity to a query.
-  • SPLiCe Cont : based on Contriever-MSMARCO (Cont) (Izacard et al., 2022), a recently proposed retrieval method that uses a transformer to rank documents based on their similarity to a query.
-Baselines One standard approach, commonly used in LLM training pipelines, is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context.
-This is known as example packing (Brown et al., 2020).
-Another popular approach restricts sampled documents to the same meta-class (for example, Wikipedia article, source code) (Groeneveld et al., 2024; Zhao et al., 2024).
-We compare SPLiCe against both.
-SPLiCe computational efficiency Given the dataset sizes used in training LLMs, computational efficiency plays a non-trivial role.
+  return $\texttt{CONCAT}(\texttt{TRIM}(\texttt{ORDER}(C),L))$
+SPLiCe Computational Efficiency
+Given the dataset sizes used in training LLMs, computational efficiency plays a crucial role.
 SPLiCeRepo is the fastest and easiest to implement but requires additional directory structure, i.e., it does not apply to general web data.
 SPLiCeBM25 uses a bag of words BM25 method that lacks deeper semantic encoding.
-SPLiCe Cont requires calculating embeddings for each document and retrieval based on the vector inner-product.
-The latter step can be done efficiently using fast approximate inner-product search, e.g., Faiss (Johnson et al., 2017).
-Burstiness SPLiCe conjecturally falls into the framework presented in Chan et al.
-(2022), which shows that the distributional properties of the training data affect the in-context capabilities of transformer models.
-In particular, it indicates the importance of "burstiness", i.e., a flatter frequency distribution with a relatively higher mass on the rare, long-tail tokens appearing in a sequence.
-In Appendix K, we show that SPLiCe increases the burstiness of the training data (measured in terms of Zipf’s coefficient of token frequency) compared to the example packing baseline.
+However, it was observed to have strong generalization properties (Thakur et al.
+2021).
+SPLiCeCont requires calculating embeddings for each document and retrieval based on the vector inner-product, but can have poorer generalization properties than BM25 (Thakur et al.
+2021).
+The retrieval step can be done efficiently using a fast approximate max IP search, e.g., Faiss (Johnson, Douze, and Jégou 2017).
+To reduce the number of times the training sample requires just copy-paste abilities and improve training step efficiency, we employ the StarCoder (Li et al.
+2023b) dataset, which was deduplicated using the pipeline from (Allal et al.
+2023).
 3 Experiments
 In this section, we show that SPLiCe improves the long context performance of large-scale language models.
-To this end, we use $3$B, $7$B, and $13$B parameter models and evaluate on tasks that test in-context learning (Section 3.2.1), question answering (Section 3.2.2), and in-context information retrieval capabilities (Section 3.2.3).
-Additionally, we provide perplexity results in Appendix E.
-Next, we show that SPLiCe can improve the core model capabilities by testing the short context performance, see Section 3.2.5.
-Finally, in Section 3.3, we tune over 30 medium ($270$M parameters) models using different data mixtures and SPLiCe parameters, to analyze its design choices.
-An important finding of our work is that presented improvements occur during a relatively short, compared to pre-training, fine-tuning.
+To this end, we use $3$B, $7$B, and $13$B parameter models.
+First, in Section 3.3, we focus on tasks that test in-context learning, question answering, and in-context information retrieval.
+Next, we show that SPLiCe can improve the core model capabilities by testing its short context performance.
+Finally, in Section 3.4, we train over 40 medium-size models ($270$M parameters) using different data mixtures and SPLiCe parameters to analyze various design choices, robustness to noise, and scaling properties.
+An important finding of our work is that presented improvements occur during a relatively short fine-tuning.
 To be more precise, $3$B models were tuned on $5.4$B tokens, whereas $7$B and $13$B models were tuned on $2$B tokens.
-The $3$B and $7$B start from OpenLLaMAv2 (Geng and Liu, 2023) that was pre-trained on $1$T tokens, the $13$B parameter models start from CodeLlama (Rozière et al., 2023) that was pre-trained on $2.5$T tokens.
+3.1 Baselines
+We consider two popular baselines used in LLM training pipelines.
+The first one is Example Packing (Brown et al.
+2020), used in the training of GPT-3 models.
+It constructs training samples by randomly sampling documents from the dataset and separating them with BOS/EOS tokens.
+The second one, which we call Within-Domain Example Packing takes random documents from the same meta-class (for example, Wikipedia, C source code) and concatenates them to create a training sample (Groeneveld et al.
+2024; Zhao et al.
+2024).
+We compare SPLiCe against both baselines.
+As we note no clear benefit of Within-Domain Example Packing over Example Packing in fine-tuning case (see Table 21 in Appendix B.3 ) in the main body of the paper we compare only against a more established Example Packing.
+We visualize the differences between baselines in Figure 2.
 3.2 Experimental Setup
-For $3$B experiments, we continue training on a 50/50 mixture of RedPajama, prepared in the standard way, and C prepared using SPLiCe BM25 .
-For $7$B and $13$B ones, we continue the training on a 50/25/25 mixture of RedPajama (50) prepared in the standard way, StackExchange (25) and C (25) prepared using SPLiCe BM25 .
-StackExchange is part of RedPajama TogetherComputer (2023), and C data come from StarCoder (Li et al., 2023b).
-Including the standard RedPajama aims to prevent the model from overfitting to artificially created long documents.
-We refer to Appendix A for a detailed description of the model architectures.
-We train with $32$K context length.
-We use a batch size of $256$K tokens per step and the learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng and Liu, 2023).
-For $13$B model we use batch of $512$K tokens per step.
-Regarding the context extensions method, we use FoT (Tworkowski et al., 2023), and the CodeLlama (CL) (Rozière et al., 2023) approach. More explicitly, we test eight models:
+For $3$B model experiments, we fine-tune on a 50/50 mixture of RedPajama, prepared in the standard way, and C prepared using SPLiCeBM25.
+For $7$B and $13$B ones, we fine-tune on a 50/25/25 mixture of RedPajama (50) prepared in the standard way, StackExchange (25) and C (25) prepared using SPLiCeBM25.
+StackExchange is part of RedPajama (TogetherComputer 2023), and C data come from StarCoder (Li et al.
+2023b).
+Including the standard RedPajama aims to prevent the model from overfitting to artificially created documents and is inspired by (Ouyang et al.
+2022; Rozière et al.
+2023).
+We analyze the impact of data mixture in Section 3.4.
+We fine-tune with $32$K context length.
+We employ the Focused Transformer (FoT) (Tworkowski et al.
+2023) and CodeLlama (CL) context extension methods (Rozière et al.
+2023).
+We use a batch size of $256$K ($512$K, resp.) tokens per step for $3$B and $7$B ($13$B, resp.) models.
+We set the learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng and Liu 2023).
+In the next section, we test eight models:
 […]
-where Baseline denotes the standard (example packing) data preparation method where context is created by sampling random documents from the corpus and separating them with BOS/EOS tokens.
-We provide results regarding the within-domain sampling baseline in Appendix C.2.
+where EP denotes the standard Example Packing (Brown et al.
+2020) method (serving as baseline) where context is created by sampling random documents from the corpus and separating them with BOS/EOS tokens (see Figure 2).
+We provide results regarding the Within-Domain Example Packing baseline in Appendix C.2.
 If not stated otherwise, in SPLiCe we use $k=1$ and the identity permutation as Order in the Algorithm 1.
@@
 3.3 Experimental Results
-3.2.1 In-Context learning
-To evaluate the improvements of in-context learning abilities, we use TREC (Li and Roth, 2002; Hovy et al., 2001) and DBpedia (Lehmann et al., 2015), which are text classification tasks.
-For TREC, to test the long context utilization, we vary the number of in-context examples in $\{780,1560\}$ (which correspond roughly to $\{16,32\}$K context length).
-We sample these examples multiple times and measure classification accuracy on the test set.
-In Table 1, we observe that SPLiCe improves the average performance across different context lengths and model sizes.
-We follow a very similar protocol for DBpedia ($\{16,32\}$K context length corresponding to $\{190,380\}$ in-context examples), confirming that SPLiCe improves the context utilization.
-In Appendix I, we study the distribution of improvements with respect to the choice of the in-context examples, showing the stochastic domination of SPLiCe over the baseline.
-Table 1: We test the average classification performance on TREC (Li and Roth, 2002; Hovy et al., 2001).
-We average across $50$ sets of in-context examples for $3$B models, $10$ sets for $7$B models, and $5$ sets for $13$B models.
-We use $\Delta$ [confidence interval] to denote the mean improvement and its $95\%$ bootstrap confidence intervals (see Appendix I).
-For DBpedia, we average results across $40$ sets of in-context examples for $3$B and $7$B models and $5$ for $13$B ones.
-Due to the size of the DBpedia evaluation dataset, for each set of in-context examples, we subsample a subset of $500$ elements of the evaluation set.
-Task TREC DBpedia
-Model Context Baseline SPLiCe $\Delta$ [conf interv] Baseline SPLiCe $\Delta$ [conf interv]
+In-Context Learning
+In this section, we ask the following research questions: Does SPLiCe improve in-context learning abilities?
+If so, with what context length is it the case?
+To answer those questions, we evaluate the accuracy of our models on TREC (Li and Roth 2002; Hovy et al.
+2001) and DBpedia (Lehmann et al.
+2015), which are text classification tasks.
+For TREC we test $\{2,16,32\}$K context lengths, which correspond to $\{90,780,1560\}$ in context examples, respectively.
+For DBpedia, we test $\{16,32\}$K context lengths, which correspond to $\{190,380\}$ in-context examples, respectively, and omit the 2K length due to its limited capacity of 20 in-context examples.
+For each context length, we average the results across several sets of in-context examples and provide average improvement of SPLiCe and its $95$% bootstrap confidence interval (improvements are calculated per set of in-context examples, see Appendix H).
+In both tasks and all considered context lengths, we note that SPLiCe significantly improves in-context learning abilities in comparison to both Example Packing and the starting checkpoint.
+We hypothesize that by increasing the amount of potentially relevant information in context, SPLiCe allows the model to learn longer and better context lookups.
+We further study this in Section 3.5, where we analyze SPLiCe using the framework from (Chan et al.
+2022).
+We present the main results in Tables 1, 2 and additional in Table 35 in Appendix O.
+Additionally in Appendix H we analyze results per-set of in-context examples and show that SPLiCe achieves stochastic domination over Example Packing.
+TREC
+Model Context EP SPLiCe $\Delta$ [conf interv]
 […]
-$16$K 68.9 76.9 8.0 [6.9,9.3]{6.9,9.3]}6.9 , 9.3 ] 79.1 82.0 2.9 [2.5,3.4]{2.5,3.4]}2.5 , 3.4 ]
+$16$K 68.9 76.9 8.0 [6.9,9.3]{6.9,9.3}]6.9 , 9.3 ]
 […]
-$16$K 74.0 79.0 5.0 [3.4,6.0]{3.4,6.0]}3.4 , 6.0 ] 83.6 85.6 2.0 [1.5,2.5]{1.5,2.5]}1.5 , 2.5 ]
+$16$K 74.0 79.0 5.0 [3.4,6.0]{3.4,6.0}]3.4 , 6.0 ]
 […]
-$16$K 81.4 82.5 1.1 [0.2,1.6]{0.2,1.6]}0.2 , 1.6 ] 96.2 96.4 0.2 [0.0,0.3]{0.0,0.3]}0.0 , 0.3 ]
+$16$K 81.4 82.5 1.1 [0.2,1.6]{0.2,1.6}]0.2 , 1.6 ]
 […]
-$16$K 88.2 91.2 3.0 [1.9,3.5]{1.9,3.5]}1.9 , 3.5 ] 95.8 96.8 1.0 [0.6,1.5]{0.6,1.5]}0.6 , 1.5 ]
-Table 2: We measure question answering over long input using Qasper (Dasigi et al., 2021) (2-shot setting) and HotPotQA (Yang et al., 2018) (10-shot setting).
-For Qasper, we use the implementation from Language Model Evaluation Harness (Gao et al., 2021).
-For HotPotQA, we average results across $7$ sets of in-context examples for $3$B and $7$B models and $2$ for $13$B ones, with $\Delta$[confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence intervals.
-Note that in the $3$B model case, despite using SPLiCe for code data only, we still have improvements in non-code tasks.
-Task QASPER HotPotQA
-Model Context Baseline SPLiCe Context Baseline SPLiCe $\Delta$ [conf interv]
+$16$K 88.2 91.2 3.0 [1.9,3.5]{1.9,3.5}]1.9 , 3.5 ]
+Table 1: We test the classification accuracy on TREC (Li and Roth 2002; Hovy et al.
+2001).
+We average across $50$ sets of in-context examples for $3$B models, $10$ for $7$B models, and $5$ for $13$B models.
+$\Delta$ [ci] denotes the mean improvement and its $95\%$ bootstrap confidence intervals (see Appendix H).
+DBpedia
+Model Context EP SPLiCe $\Delta$ [conf interv]
+[…]
+$16$K 79.1 82.0 2.9 [2.5,3.4]{2.5,3.4}]2.5 , 3.4 ]
+[…]
+$16$K 83.6 85.6 2.0 [1.5,2.5]{1.5,2.5}]1.5 , 2.5 ]
+[…]
+$16$K 96.2 96.4 0.2 [0.0,0.3]{0.0,0.3}]0.0 , 0.3 ]
+[…]
+$16$K 95.8 96.8 1.0 [0.6,1.5]{0.6,1.5}]0.6 , 1.5 ]
+Table 2: We average results across $40$ sets of in-context examples for $3$B and $7$B models and $5$ for $13$B.
+Due to the size of the DBpedia dataset, for each set of in-context examples, we sample a subset of $500$ elements of the evaluation set.
+Question Answering and In-Context Retrieval
+In this section, we ask the following research question: Does fine-tuning on SPLiCe prepared data result in improved question-answering abilities?
+To answer the question, we utilize popular long context benchmarks such as Needle In A Haystack (Kamradt 2023) and lost-in-the-middle key-value retrieval (Liu et al.
+2023), along with Qasper (Dasigi et al.
+2021) from SCROLLS (Shaham et al.
+2022), HotPotQA (Yang et al.
+2018) passkey (Mohtashami and Jaggi 2023) and parts of RULER (Hsieh et al.
+2024) tasks.
+On Needle In A Haystack, we observe that the model fine-tuned on data prepared by SPLiCe strongly outperforms the model fine-tuned on data prepared by Example Packing.
+To be more precise model fine-tuned with SPLiCe can answer the question no matter the location of the relevant piece of information.
+Whereas the model trained with Example Packing only manages to answer correctly when the information is close to the question (see Figure 1).
+We also test our models on the lost-in-the-middle key-value retrieval task (Liu et al.
+2023), and observe that SPLiCe helps on hard-to-retrieve positions (see Figure 3).
+The main difference between those two tasks is that in the lost-in-the-middle key-value retrieval task, the input is highly structured (dictionary of random $128$ bit UUIDs, see Appendix D for details) and the objective of the model is to retrieve the value assigned to a given key.
+On the other hand, in the Needle In A Haystack, a piece of information is placed inside a large coherent text, and the model is asked a question related to this information (see Appendix N for details).
+We additionally evaluate our models on Qasper (Dasigi et al.
+2021), HotPotQA (Yang et al.
+2018) passkey retrieval (Mohtashami and Jaggi 2023) and RULER (Hsieh et al.
+2024) and observe that SPLiCe results in improvements over both the Example Packing and starting checkpoint.
+We present the results in Appendix K.
+Short Context Evaluation
+One challenge in long-context fine-tuning is the degradation of short-context performance (Dubey, Jauhri, and et al.
+2024).
+This can be overcomed by upsampling the short-context data and more gradual context extension (Dubey, Jauhri, and et al.
+2024).
+We note that those approaches are compatible with SPLiCe and instead focus on comparing SPLiCe with Example Packing in a single-step context extension setup.
+We observe that SPLiCe seems to be either better or on par with Example Packing (see Table 3).
+What is intriguing is that for the $13$B parameter model, SPLiCe even improves the short context performance on GSM8K (Cobbe et al.
+2021) by ($+1.7$) over the starting checkpoint.
+We hypothesize that GSM8K is a much more attention-demanding task than MMLU (Hendrycks et al.
+2021), as it requires extracting relevant pieces of information, composing a chain of thought, and writing the final answer.
+Whereas MMLU is a well-established collection of tests spanning multiple domains.
+We hypothesize that the improvement does not occur in smaller models due to their low scores on GSM8K, as we get similar results when evaluating on code in Appendix L.
+Model MMLU GSM8K
+STEM HUM All
+$7$B ST CHKP 33.6 42.1 40.8 8.0
+$7$B 2K Tuned 33.7 40.8 39.4 8.4
 […]
@@
 […]
-3.2.2 Question answering
-We also show that SPLiCe models have improved question answering capabilities.
-To this end, we use Qasper (Dasigi et al., 2021) from SCROLLS (Shaham et al., 2022) and HotPotQA (Yang et al., 2018).
-Qasper contains questions regarding research papers provided as a part of the input context.
-In Table 2, we observe that SPLiCe improves the two-shot performance.
-Similarly, we observe improvements on HotPotQA, which tests the ability to aggregate knowledge across multiple documents.
-We stress that those results measure few-shot performance, as neither of the datasets was used during training.
-3.2.3 Lost in the middle
-The key retrieval task introduced in (Liu et al., 2023) has become a routine diagnostic tool to study context capabilities.
-Specifically, the model is presented with a list of key-value pairs, which are $128$-bit UUIDs, and is tasked to retrieve the value for a given key.
-Even though very simple, the task proved to be challenging for many open-source language models.
-Specifically, LMs tend to achieve better performance when the relevant information is either in the beginning or the end of the input context, while struggling in the middle.
-See Appendix D for the task details.
-In Figure 2, we present the key retrieval performance of $7$B models trained with CodeLlama (Rozière et al., 2023) context extension method.
-We note that despite relatively short tuning, SPLiCe significantly helps on hard-to-retrieve positions (the position of the key in the range $[25,100]$).
-We present more details in Appendix D.
-Figure 2: Key-value retrieval performance on a dictionary of $300$ key-value pairs ($\approx$ $24$K tokens).
-The $7$B CL model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions in the middle.
-The details about this task can be found in Appendix D.
-Each position was evaluated using $500$ examples.
-3.2.4 Ordering of examples
-We typically use the identity ordering as Order in Algorithm 1 to merge documents into a single context, as we found that in most cases, it performs best.
-In Appendix F, we found that random shuffling is slightly better in some cases.
-Specifically, this is the case of large $7$B CodeLlama models, see Table Table 18.
-We hypothesize that random ordering forces the model to make use of the full space of the RoPe positional encoding.
-Whereas the identity ordering, also used in (Shi et al., 2023b), skews the model toward paying more attention to fragments of text that are not too far away.
-This suggests an interesting research direction on the intersection of data preparation and positional embeddings.
-3.2.5 Short context evaluation
-An undesirable side effect of fine-tuning for extended context (e.g., with SPLiCe) could be performance degradation on shorter contexts.
-To test this, we evaluate the short context performance of our models using GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021) and TREC (Li and Roth, 2002; Hovy et al., 2001).
-We present the results in Table 3.
-We note that for smaller models, both SPLiCe and Baseline show some short context degradation, whereas the larger $13$B CodeLlama Model SPLiCe brings significant improvements in short context performance on GSM8K ($+1.7$), TREC ($+3.3$) and maintains the MMLU performance.
-We provide an additional evaluation on code in Appendix L.
-Table 3: We evaluate our models on MMLU ($5$-shot), GSM8K ($8$-shot) and TREC ($90$-shot).
-We provide an additional comparison with their starting checkpoint.
-For $7$B case, we additionally compare with a model tuned with $2$k context length on the same data.
-For each task, we highlight best results up to $1$ point.
-Model MMLU GSM8K TREC
-STEM Humanities Social Other Mean
-$3$B Starting Chkpt 24.4 25.6 26.3 27.8 25.9 4.3 58.4
+$13$B ST CHKP 36.6 48.2 44.3 21.4
 […]
 […]
-$7$B Starting Chkpt 33.6 42.1 46.0 44.5 40.8 8.0 59.6
-$7$B 2K Tuned 33.7 40.8 44.6 41.1 39.4 8.4 62.3
-[…]
-[…]
-[…]
-[…]
-$13$B Starting Chkpt 36.6 48.2 50.6 45.4 44.3 21.4 78.7
-[…]
-[…]
+Table 3: We evaluate our models on MMLU ($5$-shot), GSM8K ($8$-shot CoT).
+We provide an additional comparison with their starting checkpoint.
+For the $7$B case, we additionally compare with a model tuned with $2$k context length on the same data.
+For each task, we highlight the best results up to $1$ point.
+For $3$B model results see Appendix I.
 3.4 Detailed Study with Medium Models
-In Table 5 and Table 5, we present a comprehensive examination of the impact of document packing on long-context performance using $270$M parameter models.
-Using the smaller size allows us to train over $30$ models and, thus, more accurately estimate the impact of various design choices.
-Training protocol
-Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer, 2023).
-Subsequently, we continue training using $1$B tokens with the context extended to $32$K on a mixture of the original RedPajama data (TogetherComputer, 2023) and long context data created using SPLiCe/Baseline.
-The amount of pre-training tokens is based on scaling laws from (Hoffmann et al., 2022) and constants calculated for GPT-like models in (Karpathy, 2022).
-We employ the Focused Transformer (FoT) objective Tworkowski et al.
-(2023) for context extension (unless stated otherwise).
-This approach is motivated by practical factors, viz.
-training with short context length expedites the process, while context scaling can be achieved by finetuning on a relatively small amount of tokens, as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023).
-Loosely inspired by (Ouyang et al., 2022; Rozière et al., 2023), in the latter phase, long context data (i.e., prepared with SPLiCe) constitutes half of the mixture.
-We also check how SPLiCe works with other context extension methods in Appendix B.2, and 64K context length in Appendix J.
-Evaluation
-We measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets employing a context length of $32$K.
-The selection of these datasets is motivated by the fact that they can benefit from long-context information as demonstrated in (Chen et al., 2023; Li et al., 2023b).
-For example, functions are often re-utilized, and similar terms are employed across papers.
-We exclude documents with fewer than $32$K tokens and truncate those exceeding this length.
-In Appendix B.1 we evaluate our models using short context data confirming no performance degradation with respect to the base model.
-For information regarding the training and evaluation data, see Appendix G.
-[…]
+In Table 7, Table 7, and Table 8, we present a comprehensive examination of the impact of document packing on long-context performance using $270$M parameter models, showing that SPLiCe brings consistent improvement in long context language modeling.
+In Table 7, we scale context to 64K and observe even greater benefits over the Example Packing.
+We also expand our results to 131K and 160K context length in Tables 19 and 20 in Appendix.
+In Table 7, we show that SPLiCe is quite robust to the non-accurate retriever.
+What is more results in Table 7 clearly show that SPLiCe is an improvement over Within-Domain Example Packing.
+In particular, we note that the more noise we add, the closer SPLiCe is to Within-Domain Example Packing (semantically), and that with 100% noise SPLiCe turns into Within-Domain Example Packing (this is because we use SPLiCe to prepare data coming from a single domain).
+Training and Evaluation
+Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer 2023).
+Subsequently, we fine-tune using $1$B tokens with the context extended to $32$K on a mixture of the original RedPajama data (TogetherComputer 2023) and long context data created using SPLiCe/EP.
+We employ the Focused Transformer (FoT) (Tworkowski et al.
+2023) for context extension (unless stated otherwise).
+We measure perplexity on held-out portions of the arXiv (Azerbayev, Piotrowski, and Avigad 2022) and StarCoder (Li et al.
+2023b) datasets.
+The selection of these datasets is motivated by the fact that they can benefit from long-context information as demonstrated in (Chen et al.
+2023; Li et al.
+2023b).
+We provide additional details in Appendix M.
 Altered Method arXiv Code Code &
 Data Haskell Python CUDA All arXiv
-C SPLiCe BM25 5.46 (.09) 3.20 (.17) 2.81 (.12) 2.22 (.11) 2.94 (.13) 3.10 (.13)
-SPLiCe Cont 5.48 (.07) 3.22 (.15) 2.82 (.11) 2.23 (.10) 2.96 (.11) 3.11 (.12)
-SPLiCe Repo 5.47 (.08) 3.22 (.15) 2.83 (.10) 2.24 (.09) 2.96 (.11) 3.12 (.11)
-Baseline 5.55 3.37 2.93 2.33 3.07 3.23
 C# SPLiCeBM25 5.52 (.13) 3.33 (.25) 2.90 (.17) 2.46 (.19) 3.11 (.20) 3.26 (.20)
@@
 SPLiCeRepo 5.53 (.12) 3.35 (.23) 2.91 (.16) 2.49 (.16) 3.12 (.19) 3.27 (.19)
-Baseline 5.65 3.58 3.07 2.65 3.31 3.46
+EP 5.65 3.58 3.07 2.65 3.31 3.46
 Python SPLiCeBM25 5.47 (.10) 3.25 (.21) 2.53 (.09) 2.41 (.15) 3.02 (.15) 3.17 (.15)
@@
 SPLiCeRepo 5.48 (.09) 3.27 (.19) 2.54 (.08) 2.44 (.12) 3.03 (.14) 3.18 (.14)
-Baseline 5.57 3.46 2.62 2.56 3.17 3.32
+EP 5.57 3.46 2.62 2.56 3.17 3.32
 Wikipedia SPLiCeBM25 5.64 (.09) 3.82 (.15) 3.26 (.11) 2.87 (.13) 3.55 (.13) 3.68 (.13)
 SPLiCeCont 5.65 (.08) 3.87 (.10) 3.30 (.07) 2.92 (.08) 3.59 (.09) 3.72 (.09)
-Baseline 5.73 3.97 3.37 3.00 3.68 3.81
+EP 5.73 3.97 3.37 3.00 3.68 3.81
 StackEX SPLiCeBM25 5.07 (.07) 3.88 (.06) 3.32 (.04) 2.89 (.05) 3.60 (.05) 3.69 (.05)
 SPLiCeCont 5.09 (.05) 3.91 (.03) 3.35 (.01) 2.93 (.01) 3.63 (.02) 3.73 (.01)
-Baseline 5.14 3.94 3.36 2.94 3.65 3.74
-Table 5: Perplexity fine-tune on a $/50$ 50 data mixture of RedPajama and C code.
-We report the mean and standard deviation.
-Interestingly, training on the code data with SPLiCe improves general long-context performance on arXiv.
+EP 5.14 3.94 3.36 2.94 3.65 3.74
+[…]
 Long Context Method arXiv Code Code &
@@
 SPLiCeRepo 5.474 $\pm$ .007 2.827 $\pm$ .006 2.958 $\pm$ .009 3.115 $\pm$ .009
-Baseline 5.550 $\pm$ .002 2.931 $\pm$ .008 3.073 $\pm$ .006 3.228 $\pm$ .005
-3.4 Summary of the results
-Structuring data improves performance.
-We observe (see Tables 15) that all tested variants of SPLiCe outperform Baseline, often significantly, regardless of the training and eval dataset.
-This indicates that the structure of the data is important for long-context training.
-The positive effects transfer between domains.
-We observe, see Tables 2 and 5, that the positive effects of SPLiCe transfer between domains; when training on code and evaluating on Qasper/arXiv and when training on web data and evaluating on code.
-This is a very interesting effect, which indicates that training induces generic long-context capabilities.
-Following that, we conjecture that better data mixtures and synthetic long-context data could bring further benefits.
-SPLiCe improves the performance of various context extension methods SPLiCe only changes the input structure, and conjecturally, it should work with any context extension method.
-Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023).
-FoT extends the context only in a couple of attention layers and does not change the positional embeddings of the remaining layers.
-In Table 9 in Appendix B.2, we confirm that SPLiCe integrates seamlessly with methods that extend the context in all layers.
-Specifically, we test adjusting Rotary Positional Encodings according to either the recipe as in CodeLlama (Rozière et al., 2023), or YaRN (Peng et al., 2023), or left without changes.
-Training is stable.
-We prepared three subsets of the C code dataset and ran the tested methods on each of them.
-In Table 5, we report the mean perplexity and its standard deviation.
-We observe that the differences between the subsets are minimal, which indicates training stability and confirms the statistical significance of our results.
+EP 5.550 $\pm$ .002 2.931 $\pm$ .008 3.073 $\pm$ .006 3.228 $\pm$ .005
+Table 5: Perplexity fine-tune on a $/50$ 50 data mixture of RedPajama and C code.
+We report the mean and standard deviation.
+Interestingly, training on the code data with SPLiCe improves general long-context performance on arXiv.
+Altered Method arXiv Code Code &
+Data Haskell Python CUDA All arXiv
+C# SPLiCeBM25 4.86 (.15) 2.60 (.19) 2.66 (.16) 2.32 (.19) 2.75 (.19) 2.88 (.19)
+SPLiCeCont 4.88 (.13) 2.62 (.17) 2.67 (.15) 2.34 (.17) 2.77 (.17) 2.90 (.17)
+SPLiCeRepo 4.88 (.13) 2.62 (.17) 2.68 (.14) 2.35 (.16) 2.77 (.17) 2.90 (.17)
+EP 5.01 2.79 2.82 2.51 2.94 3.07
+[…]
+Method SPLiCe EP
+Eval Data/Noise 0% 10% 25% 50% 75% 90% -
+arXiv 5.46 5.47 5.48 5.50 5.53 5.55 5.55
+Code Python 2.81 2.82 2.83 2.86 2.89 2.92 2.93
+All 2.94 2.95 2.97 3.01 3.04 3.06 3.07
+Code & arXiv 3.10 3.11 3.13 3.16 3.19 3.22 3.23
+Table 7: We test the robustness of SPLiCe to noisy retriever.
+We achieve this by preparing data using BM25 retriever that with probability $p$ returns a random document instead of the most related one.
+We note that SPLiCe is quite robust and only with $p=0.9$ approaches Example Packing.
+Altered Method arXiv Code Code
+Data Python All arXiv
+25% SPLiCe 5.43 2.91 3.08 3.22
+25% EP 5.49 3.00 3.19 3.34
+50% SPLiCe 5.46 2.81 2.94 3.10
+50% EP 5.55 2.93 3.07 3.23
+75% SPLiCe 5.61 2.80 2.86 3.05
+75% EP 5.73 2.94 3.04 3.23
+Table 8: We note that SPLiCe beats the EP perplexity when trained with various proportions of SPLiCeBM25/EP prepared C data (the remaining data is unaltered RedPajama).
+3.5 Properties of SPLiCe Generated Data
+SPLiCe conjecturally falls into the framework presented in (Chan et al.
+2022), which shows that the distributional properties of the training data affect the in-context capabilities of transformer models.
+In particular, it indicates the importance of ”burstiness”, i.e., a flatter frequency distribution with a relatively higher mass on the rare, long-tail tokens appearing in a sequence.
+In Table 9, we show that SPLiCe increases the burstiness of the training data (measured in terms of Zipf’s coefficient of token frequency) in comparison to the Example Packing.
+Training Data Method Zipf’s Coefficient
+C SPLiCeBM25 ${1.512}_{(0.055)}$
+EP ${1.593}_{(0.025)}$
+StackEx SPLiCeBM25 ${1.643}_{(0.026)}$
+EP ${1.664}_{(0.013)}$
+Table 9: Zipf’s coefficient of token frequency on EP and SPLiCe along with standard deviation. A lower Zipf’s coefficient represents a more significant burstiness property.
 4 Related Work
 There is an increasing number of works aiming to study the role of data in LLM training in detail.
-For instance, (Levine et al., 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training examples leads to better sentence embeddings and improves open-domain question-answering performance.
-Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include intrinsic tasks.
+For instance, (Levine et al.
+2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training samples leads to better sentence embeddings and improves open-domain question-answering performance.
+Another study by (Gu et al.
+2023) introduced a pretraining framework grounded on the idea that text documents often include intrinsic tasks.
 They showed that this approach substantially boosts in-context learning.
-Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al., 2022).
-Work (Chan et al., 2022) identifies the training data’s distributional properties that affect transformer models’ in-context capabilities.
-Similarly, Han et al.
-(2023) constructs small-scale data using an iterative gradient approach and shows that such data improve in-context performance.
+Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al.
+2022).
+Work of (Chan et al.
+2022) identifies the training data’s distributional properties that affect transformer models’ in-context capabilities.
+Similarly, (Han et al.
+2023) constructs small-scale data using an iterative gradient approach and shows that such data improve in-context performance.
 Our methodology diverges from these works in several key ways.
-Firstly, we focus on the document-level context during the training phase and long context performance, as opposed to sentence-level (Levine et al., 2022) or paragraph-level (Gu et al., 2023) granularity.
-We demonstrate the efficacy of our approach in large-scale language modeling, specifically with OpenLLaMA $3$B, $7$B, and CodeLlama $13$B.
-Secondly, we construct a tree structure of related documents through BM25/Contriever-MSMARCO retrieval and linearize this structure to form long-context examples.
-This allows for greater control over example coherence compared to relying solely on natural data structures like repository-level code.
-The gradient-based method presented in Han et al.
-(2023) can be broadly associated with retrieval used in our work; however, it differs in scale and granularity.
-Concurrently to our research, (Shi et al., 2023b) showed a method for preparing the training data that closely resembles SPLiCe with default settings ($k=1$ and the identity as Order).
-The key differences are that (Shi et al., 2023b) focuses on training from scratch while we aim to achieve long context capabilities with short and cheap fine-tuning.
-Next, we use longer context lengths, up to $32$K tokens, whereas (Shi et al., 2023b) utilizes $8$K.
-Finally, we provide detailed analysis regarding the design choices, such as the benefits of data reordering in Section 3.2.4 and different values of $k$ in Appendix C.1.
+First, while prior studies have focused on sentence-level (Levine et al.
+2022) or paragraph-level (Gu et al.
+2023) granularity, we emphasize document-level context during training, specifically targeting long-context performance.
+We validate our approach in large-scale language modeling, using models such as OpenLLaMA $3$B, $7$B, and CodeLlama $13$B.
+Second, we construct a tree structure of related documents using BM25/Contriever-MSMARCO retrieval, which we then linearize to form long-context samples.
+This approach allows for greater control over the coherence of samples, compared to relying solely on natural data structures like repository-level code.
+While the gradient-based method in (Han et al.
+2023) shares similarities with our retrieval-based approach, our method scales to larger datasets and operates at a different granularity.
+Concurrently with our research, (Shi et al.
+2024) introduced a method for preparing training data that shares similarities with SPLiCe, particularly in its default settings ($k=1$ with identity as Order).
+However, while their approach focuses on training models from scratch, our work demonstrates that long-context capabilities can be effectively achieved through short and cost-efficient fine-tuning.
+In addition to this distinction, we employ significantly longer context lengths, extending above 64K tokens compared to the 8K tokens used in (Shi et al.
+2024), which allows for more comprehensive context handling.
+Furthermore, we provide an in-depth analysis of our design choices, such as the advantages of data reordering (detailed in Appendix J) and the impact of varying $k$ values (Appendix C.1).
+These analyses underline the effectiveness and flexibility of our approach.
+Our findings are especially pertinent in the context of recent research on the ”Physics of Language Models” (Allen-Zhu and Li 2024), which discusses the limitations of fine-tuning.
+Despite these limitations, we show that SPLiCe offers substantial and quantifiable improvements even with relatively brief fine-tuning, providing a practical advantage in enhancing long-context capabilities.
 5 Limitations and Future Work
-We show that structuring the training data is a viable way of improving the model’s performance.
-The presented method, SPLiCe, can viewed as a general framework for organizing the documents into training examples.
+We show that structuring the training data is a viable way of improving the model’s long-context performance.
+The presented method, SPLiCe, can be viewed as a general framework for organizing the documents into training samples.
 This opens multiple further research avenues.
-Degree of relevance In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Cont/Repo).
-However, recent results of (Tworkowski et al., 2023) show that introducing unrelated data to the context may help the model learn better representations.
-We leave the study of how the choice of retriever (in particular, the ratio of related and unrelated documents) affects performance for future work.
-In order to prepare the training data, SPLiCe uses each document exactly once (we mask out each used document for further retrieval).
-However, it is possible that allowing some high-quality documents to occur more than once may be beneficial.
-Retrieval granularity Another avenue for future work is to study the granularity of the pieces from which the training examples are constructed.
+Retrieval Granularity Another avenue for future work is to study the granularity of the pieces from which the training samples are constructed.
 In this work, we focus on the document-level granularity.
-However, it is possible to construct training examples from smaller pieces.
-Other context extension methods In most cases, our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023).
-Additionally, we tested three popular context extension methods on a medium scale (Naive, YaRN, and CodeLlama) and tuned a large $7$B model using the CodeLlama approach, preliminary confirming the applicability of SPLiCe.
-However, we leave more detailed studies with other context extension methods to future work.
-Other data sources One of the approaches to training long-context language models is to use conversational data (Li et al., 2023a).
-This is complementary to our method.
-SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, or StackExchange questions and answers) to further increase the number of long-context examples.
+However, it is possible to construct training samples from smaller pieces.
+Other Data Sources One of the approaches to training long-context language models is to use conversational data (Li et al.
+2023a), which is complementary to our method.
+SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, or StackExchange) to further increase the number of long-context samples.
 We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
@@
 We leave the study of how SPLiCe integrates with different data types for the future.
+In particular, in our studies, the datasets used were reasonably deduplicated.
+Neural Retriever In our work, we have utilized Contriever (Izacard et al.
+2022) in a zero-shot setup, using the first $512$ tokens to generate the embedding.
+On the other hand, BM25 had access to all the document content.
+Further study is required to determine whether SPLiCe can additionally significantly benefit from properly tuned neural retrievers.
+In particular, in our case, Contriever tended to produce samples consisting of fewer repositories than BM25.
+We leave this for future work.
 6 Conclusions
-In this work, we present SPLiCe, a method of constructing training examples for long-context language models.
+In this work, we present SPLiCe, a method of constructing training samples for long-context language models.
 It utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner.
@@
 We believe that our work indicates multiple interesting research directions for improving the performance of long-context language models with structured data.
-Acknowledgments and Disclosure of Funding
-We gratefully acknowledge that our research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC).
-We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH, CI TASK) for providing computer facilities and support within computational grant numbers PLG/2023/016265.
-Piotr Miłoś was supported by National Science Center Poland under the grant agreement 2019/35/O/ST6/03464.
+Ethical Statement
+Our work develops a generic technique that allows for improving context utilization in language models via low-cost fine-tuning.
+However, we note that it does not create any new threats, but only exacerbates existing ones.
+Therefore, we refer to the existing literature on the broader impact of language models, such as (Borgeaud et al.
+2022).
+Acknowledgments
+We are thankful for the TPU Research Cloud program, which was instrumental to our research by providing significant computational resources.
+Parts of the project were realized using the resources of IDEAS NCBR.
 Appendix A Architecture
-The architecture of our models is based on LLaMA [Touvron et al., 2023], and the architectural details can be found in Table 6.
-Briefly speaking, our architecture is similar to the one introduced in [Vaswani et al., 2017] with a few standard changes.
+The architecture of our models is based on LLaMA (Touvron et al.
+2023), and the architectural details can be found in Table 10.
+Briefly speaking, our architecture is similar to the one introduced in (Vaswani et al.
+2017) with a few standard changes.
 First, we use only the decoder without the encoder part.
@@
 Thirdly, we use the LLaMA FeedForward module.
-Additionally, we use Rotary Position Embedding [Su et al., 2021].
-For context extension, we use Focused Transformer (FoT)[Tworkowski et al., 2023], CodeLlama [Rozière et al., 2023] (CL) and YaRN Peng et al.
-[2023].
-Table 7 presents the details about both standard and long-context pretraining.
+Additionally, we use Rotary Position Embedding (Su et al.
+2021).
+For context extension, we use Focused Transformer (FoT)(Tworkowski et al.
+2023), CodeLlama (Rozière et al.
+2023) (CL) and YaRN (Peng et al.
+2023).
+Table 11 presents the details about both standard and long-context pretraining/fine-tuning.
 We use AdamW as an optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.95$.
-Table 6: Architecture details. Focused Transformer context extension is applied in continued pretraining for $32$K context and evaluation.
 Parameter/Model Size 270M 3B 7B 13B
@@
 FoT Context Extension Layers [6, 9] [6, 12, 18] [8, 16, 24] [12, 24, 36]
+Table 10: Architecture details. Focused Transformer context extension is applied in fine-tuning for $32$K context and evaluation.
 Each tuning experiment of 270M model was done using either a TPUv3-8 or TPUv4-8 machine and took around $10$ hours.
@@
 The $7$B and $13$B models were tuned on TPUv3-128.
-Table 7: Training details.
-We pretrain a custom $270$M parameter model and take a pretrained $3$B/$7$B/ parameter OpenLLaMAv2 model [Geng, 2023] and $13$B CodeLlama [Rozière et al., 2023] model.
-Subscript denotes that parameter was specific for a context extension method with FoT referring to [Tworkowski et al., 2023] and no-FoT to other methods (Naive, YaRN [Peng et al., 2023], CodeLlama [Rozière et al., 2023]).
 Stage Parameter/Model Size 270M 3B 7B 13B
@@
 Tokens 1B 5.4B 2B 2B
+Table 11: Training details.
+We pretrain a custom $270$M parameter model and take a pretrained $3$B/$7$B/ parameter OpenLLaMAv2 model (Geng 2023) and $13$B CodeLlama (Rozière et al.
+2023) model.
+Subscript denotes that parameter was specific for a context extension method with FoT referring to (Tworkowski et al.
+2023) and no-FoT to other methods (Naive, YaRN (Peng et al.
+2023), CodeLlama (Rozière et al.
+2023)).
 Appendix B Additional Results for $270$M Models
@@
 In Table 12, we asses the short context performance of $270$M models from Section 3.4 and compare them against their starting checkpoint.
-Table 8: $2$K context perplexity evaluation of models from Section 3.3
 Tune Dataset Method arXiv Code Code &
 Haskell Python CUDA All arXiv
-Starting Chkpt 10.80 13.40 6.00 5.00 7.17 7.40
+ST CHKPT 10.80 13.40 6.00 5.00 7.17 7.40
 Wikipedia SPLiCeBM25 10.64 13.28 6.00 5.00 7.16 7.38
 SPLiCeCont 10.60 13.24 5.99 4.99 7.14 7.36
-Baseline 10.59 13.20 5.99 5.00 7.12 7.34
+EP 10.59 13.20 5.99 5.00 7.12 7.34
 Stackexchange SPLiCeBM25 9.28 12.94 5.91 4.89 7.05 7.19
 SPLiCeCont 9.26 12.88 5.89 4.88 7.04 7.18
-Baseline 9.24 13.00 5.91 4.90 7.07 7.21
+EP 9.24 13.00 5.91 4.90 7.07 7.21
 C SPLiCeBM25 10.76 12.10 5.55 3.78 6.24 6.52
 SPLiCeCont 10.76 12.06 5.53 3.76 6.21 6.50
-Baseline 10.73 11.96 5.49 3.68 6.16 6.44
+EP 10.73 11.96 5.49 3.68 6.16 6.44
 SPLiCeRepo 10.78 12.10 5.56 3.79 6.24 6.52
@@
 SPLiCeCont 10.89 12.72 5.72 4.65 6.70 6.96
-Baseline 10.83 12.73 5.70 4.65 6.69 6.95
+EP 10.83 12.73 5.70 4.65 6.69 6.95
 SPLiCeRepo 10.91 12.77 5.74 4.68 6.72 6.98
@@
 SPLiCeCont 10.76 12.28 4.09 4.43 6.36 6.64
-Baseline 10.72 12.19 3.98 4.40 6.33 6.61
+EP 10.72 12.19 3.98 4.40 6.33 6.61
 SPLiCeRepo 10.78 12.33 4.12 4.45 6.38 6.66
+Table 12: $2$K context perplexity evaluation of models from Section 3.4
 B.2 Different Context Extension Methods
 We test SPLiCe with different context extension methods.
-In Table 9, we show that SPLiCe brings improvements also when context is extended in all layers using the naive approach, the CodeLlama [Rozière et al., 2023], and [Peng et al., 2023] RoPe [Su et al., 2021] adjustment method.
-[…]
+In Table 13, we show that SPLiCe brings improvements also when context is extended in all layers using the naive approach, the CodeLlama (Rozière et al.
+2023), and (Peng et al.
+2023) RoPe (Su et al.
+2021) adjustment method.
 RoPe Method arXiv Code Code &
@@
 SPLiCeRepo 6.25 (.08) 4.87 (.16) 3.56 (.10) 2.85 (.11) 3.74 (.11) 3.89 (.11)
-Baseline 6.33 5.03 3.66 2.96 3.85 4.00
+EP 6.33 5.03 3.66 2.96 3.85 4.00
 CL SPLiCeBM25 5.74 (.02) 4.28 (.09) 3.22 (.05) 2.53 (.05) 3.34 (.06) 3.49 (.06)
 SPLiCeRepo 5.74 (.02) 4.28 (.09) 3.22 (.05) 2.54 (.04) 3.35 (.05) 3.50 (.05)
-Baseline 5.76 4.37 3.27 2.58 3.40 3.55
+EP 5.76 4.37 3.27 2.58 3.40 3.55
 YaRN SPLiCeBM25 5.77 (.02) 4.32 (.10) 3.24 (.05) 2.55 (.06) 3.37 (.07) 3.52 (.06)
 SPLiCeRepo 5.77 (.02) 4.32 (.10) 3.24 (.05) 2.56 (.05) 3.38 (.06) 3.53 (.05)
-Baseline 5.79 4.42 3.29 2.61 3.44 3.58
+EP 5.79 4.42 3.29 2.61 3.44 3.58
+[…]
 B.3 Detailed Results
 In this section, we extend results presented in Section 3.4. Details about the number of tokens used for evaluation can be found in Table 27.
-Tables 10 and 11 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method.
+Tables 15 and 14 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method.
 Table 15 contains detailed results from training on C# and Python.
-Table 11 contains results on C averaged across three different subsets of C (for details about the construction of those subsets, see Appendix G).
-Both tables show that SPLiCe outperforms the Baseline by a significant margin.
+Table 14 contains results on C averaged across three different subsets of C (for details about the construction of those subsets, see Appendix F).
+Both tables show that SPLiCe outperforms the EP by a significant margin.
 The main advantage of SPLiCeBM25/SPLiCeCont over the SPLiCeRepo approach is that it can also be used for non-structured data.
@@
 Note that training on non-code data allows us to improve the model perplexity on the arXiv dataset in comparison to the model trained on code.
-Tables 13, 14 consider models trained with YaRN [Peng et al., 2023], CodeLlama [Rozière et al., 2023] and Naive (no adjustment to RoPE) context extension methods.
-Table 15 shows that a simple artificial extension of example length via random concatenation of documents (C source files) does not help.
-Table 10: Perplexity results comparing different ways of organizing the same data.
-All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in the mentioned ways.
-For details about training, please refer to Section 3.3.
+Tables 17, 19 consider models trained with YaRN (Peng et al.
+2023), CodeLlama (Rozière et al.
+2023) and Naive (no adjustment to RoPE) context extension methods.
+Table 21 shows that a simple artificial extension of example length via random concatenation of documents within domain (C source files) does not help.
+Altered Train Data: C
+Eval/Method SPLiCeBM25 SPLiCeCont EP SPLiCeRepo
+ArXiv 5.463 $\pm$ 0.002 5.477 $\pm$ 0.005 5.550 $\pm$ 0.002 5.474 $\pm$ 0.007
+C 2.126 $\pm$ 0.020 2.134 $\pm$ 0.020 2.173 $\pm$ 0.023 2.135 $\pm$ 0.020
+C++ 2.396 $\pm$ 0.004 2.403 $\pm$ 0.002 2.467 $\pm$ 0.001 2.403 $\pm$ 0.006
+CUDA 2.219 $\pm$ 0.002 2.235 $\pm$ 0.005 2.330 $\pm$ 0.009 2.239 $\pm$ 0.004
+C# 1.939 $\pm$ 0.000 1.948 $\pm$ 0.001 2.016 $\pm$ 0.004 1.949 $\pm$ 0.002
+Common Lisp 2.994 $\pm$ 0.072 3.042 $\pm$ 0.091 3.195 $\pm$ 0.078 3.043 $\pm$ 0.102
+Dart 2.141 $\pm$ 0.002 2.155 $\pm$ 0.004 2.268 $\pm$ 0.009 2.156 $\pm$ 0.002
+Emacs Lisp 7.857 $\pm$ 0.017 7.851 $\pm$ 0.020 8.098 $\pm$ 0.027 7.840 $\pm$ 0.019
+Erlang 2.665 $\pm$ 0.002 2.680 $\pm$ 0.003 2.769 $\pm$ 0.007 2.676 $\pm$ 0.003
+Fortran 3.306 $\pm$ 0.010 3.335 $\pm$ 0.010 3.554 $\pm$ 0.010 $3$$\pm$ 0.012
+Go 1.910 $\pm$ 0.001 1.924 $\pm$ 0.007 1.999 $\pm$ 0.002 1.924 $\pm$ 0.006
+Groovy 3.009 $\pm$ 0.001 3.026 $\pm$ 0.006 3.147 $\pm$ 0.013 3.025 $\pm$ 0.007
+Haskell 3.198 $\pm$ 0.001 3.221 $\pm$ 0.001 3.371 $\pm$ 0.008 3.220 $\pm$ 0.008
+Java 2.075 $\pm$ 0.002 2.086 $\pm$ 0.001 2.161 $\pm$ 0.006 2.085 $\pm$ 0.005
+Pascal 3.492 $\pm$ 0.026 3.496 $\pm$ 0.019 3.622 $\pm$ 0.021 3.513 $\pm$ 0.014
+Python 2.810 $\pm$ 0.002 2.824 $\pm$ 0.001 2.931 $\pm$ 0.008 2.827 $\pm$ 0.006
+Mean 3.100 $\pm$ 0.004 3.115 $\pm$ 0.006 3.228 $\pm$ 0.005 3.115 $\pm$ 0.009
+Table 14: To check the statistical significance of our results, we prepare three subsets of C (details in Appendix F) and train the models on a 50/50 mixture of RedPajama data (organized in the standard way) and C data organized using one of the methods.
+Note that the standard deviation is much lower than the perplexity improvements from using SPLiCe.
 Altered Train Data: C#
-Eval/Method SPLiCe BM25 SPLiCe Cont Baseline SPLiCe Repo
+Eval/Method SPLiCeBM25 SPLiCeCont EP SPLiCeRepo
 ArXiv 5.52 5.53 5.65 5.53
@@
 Altered Train Data: Python
-Eval/Method SPLiCe BM25 SPLiCe Cont Baseline SPLiCe Repo
+Eval/Method SPLiCeBM25 SPLiCeCont EP SPLiCeRepo
 ArXiv 5.47 5.49 5.57 5.48
@@
 Mean 3.17 3.19 3.32 3.18
-Table 11: To check the statistical significance of our results, we prepare three subsets of C (details in Appendix G and train the models on a 50/50 mixture of RedPajama data (organized in the standard way) and C data organized using one of the methods.
-Note that the standard deviation is much lower than the perplexity improvements from using SPLiCe.
-Altered Train Data: C
-Eval/Method SPLiCe BM25 SPLiCe Cont Baseline SPLiCe Repo
-ArXiv 5.463 $\pm$ 0.002 5.477 $\pm$ 0.005 5.550 $\pm$ 0.002 5.474 $\pm$ 0.007
-C 2.126 $\pm$ 0.020 2.134 $\pm$ 0.020 2.173 $\pm$ 0.023 2.135 $\pm$ 0.020
-C++ 2.396 $\pm$ 0.004 2.403 $\pm$ 0.002 2.467 $\pm$ 0.001 2.403 $\pm$ 0.006
-CUDA 2.219 $\pm$ 0.002 2.235 $\pm$ 0.005 2.330 $\pm$ 0.009 2.239 $\pm$ 0.004
-C# 1.939 $\pm$ 0.000 1.948 $\pm$ 0.001 2.016 $\pm$ 0.004 1.949 $\pm$ 0.002
-Common Lisp 2.994 $\pm$ 0.072 3.042 $\pm$ 0.091 3.195 $\pm$ 0.078 3.043 $\pm$ 0.102
-Dart 2.141 $\pm$ 0.002 2.155 $\pm$ 0.004 2.268 $\pm$ 0.009 2.156 $\pm$ 0.002
-Emacs Lisp 7.857 $\pm$ 0.017 7.851 $\pm$ 0.020 8.098 $\pm$ 0.027 7.840 $\pm$ 0.019
-Erlang 2.665 $\pm$ 0.002 2.680 $\pm$ 0.003 2.769 $\pm$ 0.007 2.676 $\pm$ 0.003
-Fortran 3.306 $\pm$ 0.010 3.335 $\pm$ 0.010 3.554 $\pm$ 0.010 $3$$\pm$ 0.012
-Go 1.910 $\pm$ 0.001 1.924 $\pm$ 0.007 1.999 $\pm$ 0.002 1.924 $\pm$ 0.006
-Groovy 3.009 $\pm$ 0.001 3.026 $\pm$ 0.006 3.147 $\pm$ 0.013 3.025 $\pm$ 0.007
-Haskell 3.198 $\pm$ 0.001 3.221 $\pm$ 0.001 3.371 $\pm$ 0.008 3.220 $\pm$ 0.008
-Java 2.075 $\pm$ 0.002 2.086 $\pm$ 0.001 2.161 $\pm$ 0.006 2.085 $\pm$ 0.005
-Pascal 3.492 $\pm$ 0.026 3.496 $\pm$ 0.019 3.622 $\pm$ 0.021 3.513 $\pm$ 0.014
-Python 2.810 $\pm$ 0.002 2.824 $\pm$ 0.001 2.931 $\pm$ 0.008 2.827 $\pm$ 0.006
-Mean 3.100 $\pm$ 0.004 3.115 $\pm$ 0.006 3.228 $\pm$ 0.005 3.115 $\pm$ 0.009
 Table 15: Perplexity results comparing different ways of organizing the same data.
-All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and other data organized using one of the methods.
+All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in the mentioned ways.
 For details about training, please refer to Section 3.4.
-Note that the model trained with SPLiCe on StackExchange outperforms the one trained on code on arXiv evaluation, showing the benefits of SPLiCe’s applicability to non-code data.
 Altered Train Data: StackExchange
-Eval/Method SPLiCe BM25 SPLiCe Cont Baseline
+Eval/Method SPLiCeBM25 SPLiCeCont EP
 ArXiv 5.07 5.09 5.14
@@
 Altered Train Data: Wikipedia
-Eval/Method SPLiCe BM25 SPLiCe Cont Baseline
+Eval/Method SPLiCeBM25 SPLiCeCont EP
 ArXiv 5.64 5.65 5.73
@@
 Mean 3.68 3.72 3.81
-Table 13: Perplexity results comparing different ways of organizing the same data for non-FoT models.
-All runs started from the same $270M$ model with $2048$ context and were trained for $16$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C# code is organized in one of three ways.
+Table 16: Perplexity results comparing different ways of organizing the same data.
+All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and other data organized using one of the methods.
+For details about training, please refer to Section 3.4.
+Note that the model trained with SPLiCe on StackExchange outperforms the one trained on code on arXiv evaluation, showing the benefits of SPLiCe’s applicability to non-code data.
 Altered Train Data: C#
 Context $16$K: CodeLlama
-Eval/Method SPLiCe BM25 Baseline SPLiCe Repo
+Eval/Method SPLiCeBM25 EP SPLiCeRepo
 ArXiv 5.74 5.76 5.74
@@
 Context $16$K: YaRN
-Eval/Method SPLiCe BM25 Baseline SPLiCe Repo
+Eval/Method SPLiCeBM25 EP SPLiCeRepo
 ArXiv 5.77 5.79 5.77
@@
 Table 17: Perplexity results comparing different ways of organizing the same data for non-FoT models.
-All runs started from the same $270M$ model with $2048$ context and were trained for $16$K context on a 50/50 mixture of RedPajama with a Naive context extension method.
+All runs started from the same $270M$ model with $2048$ context and were trained for $16$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C# code is organized in one of three ways.
 Altered Train Data: C#
 Context $16$K: Naive
-Eval/Method SPLiCe BM25 Baseline SPLiCe Repo
+Eval/Method SPLiCeBM25 EP SPLiCeRepo
 ArXiv 6.25 6.33 6.25
@@
 Mean 3.88 4.00 3.89
-Table 15: Perplexity results comparing different ways of organizing the same data.
-All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C code is organized in one of three ways.
+Table 18: Perplexity results comparing different ways of organizing the same data for non-FoT models.
+All runs started from the same $270M$ model with $2048$ context and were trained for $16$K context on a 50/50 mixture of RedPajama with a Naive context extension method.
+SPLiCeBM25 SPLiCeCont Baseline
+Eval/Method
+ArXiv 4.55 4.56 4.71
+C 2.02 2.03 2.12
+C++ 1.99 1.99 2.13
+CUDA 1.97 1.98 2.14
+C# 1.58 1.58 1.66
+Common Lisp 2.83 2.84 3.08
+Dart 2.13 2.14 2.28
+Emacs Lisp 5.90 5.92 6.26
+Erlang 1.73 1.74 1.83
+Fortran 2.75 2.76 3.06
+Go 1.76 1.76 1.87
+Groovy 2.32 2.32 2.41
+Haskell 2.13 2.14 2.28
+Java 1.68 1.69 1.81
+Pascal 3.82 3.81 4.05
+Python 2.49 2.51 2.63
+Mean 2.60 2.61 2.77
+Table 19: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C# code with longer $131$K context.
+We note that SPLiCe still brings significant benefits here, in particular when compared with $32$K and $64$K setups.
+SPLiCeBM25 SPLiCeCont Baseline
+Eval/Method
+ArXiv 4.36 4.36 4.51
+Python 2.73 2.74 2.85
+Mean 3.54 3.55 3.68
+Table 20: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C# code with longer $131$K context and evaluating on $160$K context.
+We note that FoT uses positional encodings that allow for such extrapolation.
 Altered Train Data: C
-Eval/Method SPLiCe BM25 Baseline DomRnd
+Eval/Method SPLiCeBM25 EP WithinDomEP
 ArXiv 5.46 5.55 5.55
@@
 Mean 3.10 3.23 3.23
+Table 21: Perplexity results comparing different ways of organizing the same data.
+All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C code is organized in one of three ways.
+SPLiCeBM25 SPLiCeCont Baseline Repo
+Eval/Method
+ArXiv 4.86 4.88 5.01 4.88
+C 2.17 2.19 2.28 2.19
+C++ 2.29 2.30 2.43 2.30
+CUDA 2.32 2.34 2.51 2.35
+C# 1.72 1.73 1.81 1.73
+Common Lisp 2.90 2.95 3.15 2.96
+Dart 2.07 2.09 2.22 2.09
+Emacs Lisp 7.40 7.43 7.84 7.42
+Erlang 2.10 2.11 2.23 2.11
+Fortran 3.15 3.19 3.50 3.20
+Go 1.82 1.83 1.92 1.83
+Groovy 2.54 2.59 2.76 2.58
+Haskell 2.60 2.62 2.79 2.62
+Java 1.83 1.85 1.96 1.84
+Pascal 3.62 3.64 3.82 3.62
+Python 2.66 2.67 2.82 2.68
+Mean 2.88 2.90 3.07 2.90
+[…]
 Appendix C Ablations
@@
 Second, how the documents are ordered.
-Table 16 indicates that $k=1$ is the best choice, though the differences are rather small.
-We found that changing the order of documents in training examples hardly matters for $270$M models.
+Table 23 indicates that $k=1$ is the best choice for naturally ocurring long documents whereas Table 24 shows that in case of RAG-style input greater values of $k$ achieve better performance, though the differences are rather small.
+We found that changing the order of documents in training samples hardly matters for $270$M models.
 We use ’standard’, as ordered by Algorithm 1, the reversed order, and random shuffling.
-Table 16: Ablation of SPLiCe hyper-parameters.
-For each ablation, we have trained the same $270M$ parameter model using different data organization methods.
-Top-$k$ corresponds to the number of descendants chosen in the $\texttt{RETRIEVE}(d,k)$ step of the Algorithm 1.
-Reverse and shuffle correspond to the final order of examples $C$ returned by the Algorithm 1 (reverse – the order of documents in $C$ is reversed, shuffle – examples are shuffled.
 Code
@@
 shuffle 2.40 3.24 2.84 2.25 2.96
+Table 23: Ablation of SPLiCe hyper-parameters.
+For each ablation, we have trained the same $270M$ parameter model using different data organization methods.
+Top-$k$ corresponds to the number of descendants chosen in the $\texttt{RETRIEVE}(d,k)$ step of the Algorithm 1.
+Reverse and shuffle correspond to the final order of examples $C$ returned by the Algorithm 1 (reverse – the order of documents in $C$ is reversed, shuffle – examples are shuffled.
+Method Code Code &
+arXiv C RAG-style Python Code arXiv
+SPLiCeBM25 k=1 5.46 3.20 2.81 2.94 3.10
+SPLiCeBM25 k=16 5.49 3.18 2.85 2.98 3.14
+Table 24: Ablation of SPLiCe hyper-parameter $k$, we can observe that increasing $k$ can bring benefits on RAG-style input data (input created by retrieving related documents with SPLiCe $k=16$) at a cost of performance on natural long documents.
+Models were tuned on a 50/50 mixture of C prepared with SPLiCe and RedPajama
 We also evaluated the influence of BOS and EOS tokens on the performance of trained models.
-As both Repo and SPLiCe methods concatenate documents to create training examples, they effectively slightly decrease the number of separating tokens compared to the Baseline.
+As both Repo and SPLiCe methods concatenate documents to create training samples, they effectively slightly decrease the number of separating tokens compared to the EP.
 However, in Appendix C.2 we included experiments showing that this has no impact on model performance.
@@
 We also evaluate the influence of BOS and EOS tokens on the performance.
-To be more precise, in all setups, training examples are separated by BOS and EOS tokens.
-As SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens.
+To be more precise, in all setups, training samples are separated by BOS and EOS tokens.
+As SPLiCe methods concatenate documents to create training samples, they effectively increase the average example length and decrease the number of separating tokens.
 To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens and concatenation of documents from the same domain, we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the C dataset (sampling without replacement).
 The results are shown in Table 25.
-We note that the difference between the Baseline and the random concatenation approach is small, and the random concatenation approach does not result in significant perplexity gains.
+We note that the difference between the EP and the random concatenation approach is small, and the random concatenation approach does not result in significant perplexity gains.
 When concatenating the documents, we concatenate them within the domain (C language in this case).
+Altered Method arXiv Code Code &
+Data Python All arXiv
+C WithinDomEP 5.554 $\pm$ 0.004 2.931 $\pm$ 0.003 3.076 $\pm$ 0.005 3.231 $\pm$ 0.005
+EP 5.550 $\pm$ 0.002 2.931 $\pm$ 0.008 3.073 $\pm$ 0.006 3.228 $\pm$ 0.005
 Table 25: Perplexity evaluation of two methods of organizing the data.
-Baseline – example packing, document equals training example.
-DomRnd – concatenate documents within the domain into examples of length bounded by $120$K characters.
-Training examples are then fed into the model and separated by BOS and EOS tokens.
-The difference is negligible, which suggests that the extension of example length in DomRnd does not help the model to utilize extended context.
+EP – example packing, document equals training sample.
+WithinDomEP – concatenate documents within the domain into examples of length bounded by $120$K characters.
+Training samples are then fed into the model and separated by BOS and EOS tokens.
+The difference is negligible, which suggests that the extension of example length in WithinDomEP does not help the model to utilize extended context.
 Experiments were performed using the $270$M parameter model.
@@
 The training data is a 50/50 mixture of RedPajama, organized in a standard way, and C data, organized using a method of choice.
-Altered Method arXiv Code Code &
-Data Python All arXiv
-C DomRnd 5.554 $\pm$ 0.004 2.931 $\pm$ 0.003 3.076 $\pm$ 0.005 3.231 $\pm$ 0.005
-Baseline 5.550 $\pm$ 0.002 2.931 $\pm$ 0.008 3.073 $\pm$ 0.006 3.228 $\pm$ 0.005
 Appendix D Key-Value Retrieval Task
-Figure 2 shows how training on SPLiCe organized data improves the performance on the key-value retrieval task proposed in [Liu et al., 2023].
+Figure 3 shows how training on SPLiCe organized data improves the performance on the key-value retrieval task proposed in (Liu et al.
+2023).
 This is a zero-shot task in which a model is prompted with a JSON dictionary and asked to retrieve a value corresponding to a specified key.
 The structure of the input is showcased below.
-Extract the value corresponding to the specified key in the JSON object below.
-JSON data: {"4ef217b7-6bc0-48c6-af35-2765f1e730f3": "068192b7-16b1-40e0-8495-61c63f979d50",  "cd6b8bdc-bc6c-4490-acb4-bc187a2dccba": "7364a26e-289f-4968-93d3-b273e882bdee",  "7d057372-4ab8-4811-8110-658c3f19fff4": "3ad075c5-b567-4201-85a7-cb31a0c91540",  "c62e192d-45e6-4646-bb88-1529c73256c9": "f0411644-1f6d-42a6-8af8-f06da66efc77",  "06134e93-e158-490e-a66c-8e3b98e12735": "50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",  "3286f978-4270-4b54-8bfa-540d7e0772e6": "075cc716-1836-4f90-9be3-53e3d4ec6585",  "4701aa05-c523-4b89-9700-64ab9c37c537": "49d86354-74c4-4256-9b3a-35e6e2b80d00",  "c8895805-e574-4f13-9fe5-89da1d8c4748": "cc91af7f-8509-4bdc-bad7-2646af68e6d2"}  "4701aa05-c523-4b89-9700-64ab9c37c537":
+Extract␣the␣value␣corresponding␣to␣the␣specified␣key␣in␣the␣JSON␣object␣below.
+JSON␣data:
+{"4ef217b7-6bc0-48c6-af35-2765f1e730f3":␣"068192b7-16b1-40e0-8495-61c63f979d50",
+␣"cd6b8bdc-bc6c-4490-acb4-bc187a2dccba":␣"7364a26e-289f-4968-93d3-b273e882bdee",
+␣"7d057372-4ab8-4811-8110-658c3f19fff4":␣"3ad075c5-b567-4201-85a7-cb31a0c91540",
+␣"c62e192d-45e6-4646-bb88-1529c73256c9":␣"f0411644-1f6d-42a6-8af8-f06da66efc77",
+␣"06134e93-e158-490e-a66c-8e3b98e12735":␣"50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",
+␣"3286f978-4270-4b54-8bfa-540d7e0772e6":␣"075cc716-1836-4f90-9be3-53e3d4ec6585",
+␣"4701aa05-c523-4b89-9700-64ab9c37c537":␣"49d86354-74c4-4256-9b3a-35e6e2b80d00",
+␣"c8895805-e574-4f13-9fe5-89da1d8c4748":␣"cc91af7f-8509-4bdc-bad7-2646af68e6d2"}
+␣"4701aa05-c523-4b89-9700-64ab9c37c537":
 We noted that FoT-trained models struggle with this task.
 This is probably due to the fact that they extend context only in a couple of layers, and the key-value retrieval task requires looking up and extracting a long sequence of letters and digits.
-Because of that, we evaluate FoT models with shorter dictionaries consisting of $75$ key-value pairs (around $6$K tokens) and show the results in Figures 3, and 3.
+Because of that, we evaluate FoT models with shorter dictionaries consisting of $75$ key-value pairs (around $6$K tokens) and show the results in Figures 4, and 4.
 For comparison, we also evaluate the $7$B CL model with this context length and show the results in Figure 4.
-Figure 3: Performance on a smaller version of key-value retrieval task from [Liu et al., 2023].
+Figure 4: Performance on a smaller version of key-value retrieval task from (Liu et al.
+2023).
 We note that FoT models (a), (b) generally struggle to retrieve tokens that are only visible to a subset of layers with extended context.
-For comparison, we show the results with a model that has extended context in all layers (c) using CodeLlama [Rozière et al., 2023] method of context extension.
+For comparison, we show the results with a model that has extended context in all layers (c) using CodeLlama (Rozière et al.
+2023) method of context extension.
 Each position was evaluated using $500$ examples.
 Appendix E Perplexity Improvements
-In Figure 4 we present perplexity improvements of $3$B FoT SPLiCe BM25 over Baseline.
+In Figure 5 we present perplexity improvements of $3$B FoT SPLiCeBM25 over EP.
 Figure 6 shows the evolution of SPLiCe model perplexity during the training.
-We follow [Anthropic, 2023] and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets.
+We follow (Anthropic 2023) and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets.
 We average perplexity across arXiv, CUDA, Haskell, and CommonCrawl datasets.
-Figure 4: Perplexity improvement with SPLiCe against the Baseline of the final models (after $21$k training steps).
+Figure 5: Perplexity improvement with SPLiCe against the EP of the final models (after $21$k training steps).
 We bucket tokens by their positions in the document and calculate the average.
-Each dot is the difference of the averages of the SPLiCe and Baseline models.
+Each dot is the difference of the averages of the SPLiCe and EP models.
 We observe that SPLiCe has smaller perplexity, and the improvements tend to be larger for tokens further in the document.
 Figure 6: Evolution of the perplexity with SPLiCe, as the model is trained on more tokens.
-See 4 for the difference with the baseline.
+See 5 for the difference with the baseline.
 As expected, SPLiCe significantly improves perplexity for tokens whose positions are very distant in the sequence.
 Perplexity for more distant tokens improves more significantly compared to tokens in the beginning, early in the training.
-Appendix F Shuffling
-Table 18: Average classification performance on TREC [Li and Roth, 2002, Hovy et al., 2001].
-We compare $7$B CL model trained on SPLiCe prepared data with different approaches to ordering the examples (different function Order in 1).
-SPLiCe-shuf denotes the model trained on data that shuffled the documents randomly in context, for SPLiCe-no-shuf Order was the identity function.
-We use $\pm$ to denote the standard deviation.
-We decided to stick with the model trained with random shuffling as it has slightly better long-context performance and lower standard deviation.
-TREC
-Model
-Context length
-(# of examples)
-SPLiCe-no-shuf SPLiCe-shuf
-$7$B CL $32$K (1560) 76.0 $\pm$2.5 76.6 $\pm$1.9
-$8$K (380) 77.1 $\pm$2.4 76.5 $\pm$1.8
-Appendix G Data preparation
-G.1 Evaluation data
+For more see Table 26.
+Steps 4K 8K 16K
+TREC 64.2 $\pm$ 5.3 72.9 $\pm$ 4.1 72.9 $\pm$ 4.1
+Table 26: Evaluation of the 3B SPLiCe model on 780-shot TREC using 10 seeds and different checkpoints. See Figure 6 for perlexity.
+Appendix F Data Preparation
+F.1 Evaluation Data
 We have taken a random subset of arXiv from Proof-pile.
@@
 Table 27 shows the number of tokens over which the perplexity was calculated.
-Table 19: Number of evaluation tokens in each of the considered datasets. For each context length $c$, we consider only documents that have not less than $c$ tokens and extract the $c$ tokens prefix.
 Eval/Method 16K 32K 64K
@@
 Python 16M 16M 16M
-G.2 Train data
+Table 27: Number of evaluation tokens in each of the considered datasets. For each context length $c$, we consider only documents that have not less than $c$ tokens and extract the $c$ tokens prefix.
+F.2 Train Data
 The StackExchange data was taken from the Proof-pile.
-To prepare the code train data, we take the StarCoder train splits mentioned in Section G.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs.
+To prepare the code train data, we take the StarCoder train splits mentioned in Section F.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs.
 We also split repos larger than $25$MB and filter out files that are longer than $30$k characters.
@@
 For natural language datasets, we extract a random subset of documents.
-Appendix H Faiss parameters
-Our experiments with SPLiCe Cont utilize Faiss [Johnson et al., 2019] for fast approximate inner-product search.
-To be more precise, we use the "IVF8192,Flat" index that we train on 262144 examples coming from the dataset.
-Appendix I Detailed accuracy improvements
+Appendix G Faiss Parameters
+Our experiments with SPLiCeCont utilize Faiss (Johnson, Douze, and Jégou 2019) for fast approximate inner-product search.
+To be more precise, we use the ”IVF8192,Flat” index that we train on 262144 examples coming from the dataset.
+Appendix H Detailed Accuracy Improvements
 The performance of in-context learning depends much on the choice of the in-context examples. To study this in more detail we study the following random variable
-$\Delta(c)=\text{ACC}_{\textsc{{SPLiCe}}}(c)-\text{ACC}_{\textsc{Baseline}}(c),$
-where $\text{ACC}_{\textsc{{SPLiCe}}}(c),\text{ACC}_{\textsc{Baseline}}(c)$ are the accuracies of the model trained with SPLiCe and Baseline respectively on the random choice of in-context examples $c$.
+$\Delta(c)=\text{ACC}_{\textsc{{SPLiCe}}}(c)-\text{ACC}_{\textsc{EP}}(c),$
+where $\text{ACC}_{\textsc{{SPLiCe}}}(c),\text{ACC}_{\textsc{EP}}(c)$ are the accuracies of the model trained with SPLiCe and EP respectively on the random choice of in-context examples $c$.
 Below we report the histograms of $\delta(c)$.
@@
 (d) Examples: $1560$, Context $32$K
-Figure 6: Histograms of accuracy improvement of SPLiCe BM25 over Baseline on TREC question classification task.
-The results are obtained by comparing the accuracy on the test set of TREC of the $3$B FoT model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples.
+Figure 7: Histograms of accuracy improvement of SPLiCeBM25 over EP on TREC question classification task.
+The results are obtained by comparing the accuracy on the test set of TREC of the $3$B FoT model trained with SPLiCe to the model trained with default data preparation method (EP) across $50$ sets of in-context examples.
 Each set of in-context examples consists of elements randomly sampled (without replacement) from the training subset of TREC.
-Note that the model trained with SPLiCe is almost always better than the Baseline.
+Note that the model trained with SPLiCe is almost always better than the EP.
 (a) Examples: $190$, Context $16$K, Model: $3$B FoT
@@
 (f) Examples: $380$, Context $32$K, Model: $7$B CL
-Figure 7: Histograms of accuracy improvement of SPLiCe BM25 over Baseline on DBPedia.
+Figure 8: Histograms of accuracy improvement of SPLiCeBM25 over EP on DBPedia.
 We sample $40$ sets of in-context examples and, for each set of in-context examples, evaluate on a random $500$ element subset of the DBPedia test set.
-Appendix J Longer context
-In Table 20, we present the perplexity results for $270$M parameter models trained as described in Section 3.3 but with 2x larger context extension – context extended from $2$K to $64$K, instead of from $2$K to $32$K.
+Appendix I Results of $3$B Models on MMLU and GSM8K
+We present the missing results in Table 28.
+Model MMLU GSM8K
+STEM HUM All
+$3$B ST CHKPT 24.4 25.6 25.9 4.3
 […]
-Altered Method arXiv Code Code &
-Data Haskell Python CUDA All arXiv
-C# SPLiCe BM25 4.86 (.15) 2.60 (.19) 2.66 (.16) 2.32 (.19) 2.75 (.19) 2.88 (.19)
-SPLiCe Cont 4.88 (.13) 2.62 (.17) 2.67 (.15) 2.34 (.17) 2.77 (.17) 2.90 (.17)
-SPLiCe Repo 4.88 (.13) 2.62 (.17) 2.68 (.14) 2.35 (.16) 2.77 (.17) 2.90 (.17)
-Baseline 5.01 2.79 2.82 2.51 2.94 3.07
-Appendix K Data distribution property
-Chan et al.
-[2022], Han et al.
-[2023] shows that the "burstiness" property of training data benefits a model’s generalisation.
-Specifically, "burstiness" presents a flatter frequency distribution, with a relatively higher mass on the rare, long-tail tokens appearing in a sequence, which results in a lower Zipf’s coefficient of token frequency.
-We also observe a burstiness property of data in SPLiCe, as shown in Table 21.
-Table 21: Zipf’s coefficient of token frequency on Baseline and SPLiCe. A lower Zipf’s coefficient represents a more significant burstiness property.
-Training Data Context Length Method Zipf’s Coefficient
-C 2K SPLiCe BM25 ${1.667}_{(0.139)}$
-Baseline ${1.717}_{(0.152)}$
-StackEx 2K SPLiCe BM25 ${1.836}_{(0.088)}$
-Baseline ${1.865}_{(0.074)}$
-C 32K SPLiCe BM25 ${1.512}_{(0.055)}$
-Baseline ${1.593}_{(0.025)}$
-StackEx 32K SPLiCe BM25 ${1.643}_{(0.026)}$
-Baseline ${1.664}_{(0.013)}$
+[…]
+Table 28: We evaluate our models on MMLU ($5$-shot), GSM8K ($8$-shot CoT).
+We provide an additional comparison with their starting checkpoint.
+In the main paper we note that the results of $3$B parameter model starting checkpoint are close to random and were moved this Appendix.
+See Table 3 for results regarding larger models.
+Appendix J Ordering of Examples
+We typically use the identity ordering as Order in Algorithm 1 to merge documents into a single context, as we found that in most cases, it performs best.
+We found that random shuffling is slightly better in some cases.
+Specifically, this is the case of large $7$B CodeLlama models, see Table Table 29.
+We hypothesize that random ordering forces the model to make use of the full space of the RoPe positional encoding.
+Whereas the identity ordering, also used in (Shi et al.
+2024), skews the model toward paying more attention to fragments of text that are not too far away.
+This suggests an interesting research direction on the intersection of data preparation and positional embeddings.
+TREC
+Model
+Context length
+(# of examples)
+SPLiCe-no-shuf SPLiCe-shuf
+$7$B CL $32$K (1560) 76.0 $\pm$2.5 76.6 $\pm$1.9
+$8$K (380) 77.1 $\pm$2.4 76.5 $\pm$1.8
+Table 29: Average classification performance on TREC (Li and Roth 2002; Hovy et al.
+2001).
+We compare $7$B CL model trained on SPLiCe prepared data with different approaches to ordering the examples (different function Order in 1).
+SPLiCe-shuf denotes the model trained on data that shuffled the documents randomly in context, for SPLiCe-no-shuf Order was the identity function.
+We use $\pm$ to denote the standard deviation.
+We decided to stick with the model trained with random shuffling as it has slightly better long-context performance and lower standard deviation.
+Appendix K HotPotQA and Qasper Results
+We evaluate our models on HotPotQA (Yang et al. 2018) and Qasper (Dasigi et al. 2021) from SCROLLS (Shaham et al. 2022) on both long (see Table 33) and short context (see Table 31 and Table 32)
+Task QASPER HotPotQA
+Model Context EP SPLiCe Context EP SPLiCe $\Delta$ [conf interv]
+[…]
+[…]
+[…]
+[…]
+Table 30: We measure question answering over long input using Qasper (Dasigi et al.
+2021) (2-shot setting) and HotPotQA (Yang et al.
+2018) (10-shot setting).
+For Qasper, we use the implementation from Language Model Evaluation Harness (Gao et al.
+2021).
+For HotPotQA, we average results across $7$ sets of in-context examples for $3$B and $7$B models and $2$ for $13$B ones, with $\Delta$[confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence intervals.
+We use F1 score for evaluation.
+Note that in the $3$B model case, despite using SPLiCe for code data only, we still have improvements in non-code tasks.
+QASPER
+Model Context Method Result
+3B FoT 32K SPLiCe 22.8
+2K SPLiCe 18.4
+2K EP 18.4
+32K ST CHKPT 6.4
+2K ST CHKPT 18.4
+7B FoT 32K SPLiCe 23.1
+2K SPLiCe 18.0
+2K EP 18.2
+2K ST CHKPT 18.3
+7B CL 32K SPLiCe 29.7
+2K SPLiCe 18.6
+2K EP 18.4
+2K ST CHKPT 18.3
+13B FoT 32K SPLiCe 32.0
+2K SPLiCe 18.9
+2K EP 19.1
+2K ST CHKPT 19.9
+Table 31: Comparison with starting checkpoint on Qasper. We note that the OpenLLaMA models were trained with context length 2K.
+HotPotQA
+Size Context Method Result
+3B FoT 32K SPLiCe 29.6
+2K SPLiCe 26.4
+2K EP 27.2
+2K ST CHKPT 27.3
+7B FoT 32K SPLiCe 27.3
+2K SPLiCe 21.0
+2K EP 19.7
+2K ST CHKPT 27.7
+7B CL 32K SPLiCe 31.2
+2K SPLiCe 24.2
+2K EP 24.3
+2K ST CHKPT 27.7
+13B FoT 2K SPLiCe 36.7
+2K SPLiCe 27.6
+2K EP 25.6
+2K ST CHKPT 32.9
+Table 32: Comparison with starting checkpoint on HotPotQA (0-shot - 2k context and 10-shot 20k context). We note that the OpenLLaMA models were trained with context length 2K.
+Task Landmark Attention RULER
+Model Method Passkey Retrieval Variable Tracking
+[…]
+EP 68.6 29.8
+[…]
 Appendix L HumanEval
-Table 22: We additionally evaluate our $7$B and $13$B parameter models on HumanEval [Chen et al., 2021].
-For $7$B models, all methods experience a decrease in performance, but SPLiCe is better at maintaining short context performance.
-For $13$B, both SPLiCe and Baseline increase the performance.
-We note that in our evaluation we do not use special tokens designed for CodeLlama $13$B code completion.
-Instead, evaluate all models in the same way.
-Because of the different evaluation pipeline our results for Starting Checkpoint are not the same as the ones presented in [Rozière et al., 2023].
+We perform an additional evaluation using HumanEval (Chen et al.
+2021) and present the results in Table 34.
+We note that similarly to GSM8K results from Table 3 SPLiCe improves the short context performance of the larger model.
+HumanEval
 Model Method Pass@1
@@
 […]
-Baseline 9.1
+EP 9.1
 […]
-Baseline 8.5
+EP 8.5
 $13$B Starting Checkpoint 23.8
 […]
-Baseline 26.8
+EP 26.8
+Table 34: We additionally evaluate our $7$B and $13$B parameter models on HumanEval (Chen et al.
+2021).
+For $7$B models, all methods experience a decrease in performance, but SPLiCe is better at maintaining short context performance.
+For $13$B, both SPLiCe and EP increase the performance.
+We note that in our evaluation we do not use special tokens designed for CodeLlama $13$B code completion.
+Instead, evaluate all models in the same way.
+Because of the different evaluation pipeline our results for Starting Checkpoint are not the same as the ones presented in (Rozière et al.
+2023).
+Appendix M $270$M Models Training, Fine-Tuning and Evaluation Details
+Training protocol
+Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer 2023).
+Subsequently, we fine-tune using $1$B tokens with the context extended to $32$K on a mixture of the original RedPajama data (TogetherComputer 2023) and long context data created using SPLiCe/EP.
+The amount of pre-training tokens is based on scaling laws from (Hoffmann et al.
+2022) and constants for GPT-like models in (Karpathy 2022).
+Fine-tuning protocol
+We employ the Focused Transformer (FoT) (Tworkowski et al.
+2023) for context extension method (unless stated otherwise).
+This approach is motivated by practical factors, viz.
+training with short context length expedites the process, while context scaling can be achieved by finetuning on a relatively small amount of tokens, as demonstrated by (Chen et al.
+2023; Tworkowski et al.
+2023).
+Loosely inspired by (Ouyang et al.
+2022; Rozière et al.
+2023), in the latter phase, long context data (i.e., prepared with SPLiCe) constitutes half of the mixture.
+We also check how SPLiCe works with other context extension methods in Appendix B.2.
+Evaluation
+We measure perplexity on held-out portions of the arXiv (Azerbayev, Piotrowski, and Avigad 2022) and StarCoder (Li et al.
+2023b) datasets employing a context length of $32$K.
+The selection of these datasets is motivated by the fact that they can benefit from long-context information as demonstrated in (Chen et al.
+2023; Li et al.
+2023b).
+For example, functions are often re-utilized, and similar terms are employed across papers.
+We exclude documents with fewer than $32$K tokens and truncate those exceeding this length.
+In Appendix B.1 we evaluate our models using short context data confirming no performance degradation with respect to the base model.
+For information regarding the training and evaluation data, see Appendix F.
+Appendix N Needle In A Haystack
+We utilize the following prompt for evaluating $7$B parameter CL models on Needle In A Haystack (Kamradt 2023).
+And the default needle that was used by the authors of the benchmark to evaluate models such as Anthropic’s Claude.
+As a context document, we utilized the “PaulGrahamEssays” option from the Needle In A Haystack evaluator.
+We use recall of keywords from the ground truth answer as the evaluation metric.
+You␣are␣given␣the␣following␣Question␣and␣Context.␣Answer␣the␣Question␣using␣the␣information␣hidden␣inside␣the␣Context.
+Question:␣{retrieval_question}
+Context:␣{context}
+Question:␣{retrieval_question}
+Answer:
+Appendix O TREC 2K Evaluation
+TREC Context 2K (90-shot)
+Model ST CHKP EP SPLiCe $\Delta$ [conf interv]
+[…]
+[…]
+[…]
+[…]
+Table 35: We assess the short-context (2K) performance on TREC (Li and Roth 2002; Hovy et al.
+2001).
+We average across $10$ sets of in-context examples and use $\Delta$ [conf interv] to denote the mean improvement of SPLiCe over the starting checkpoint and its $95\%$ bootstrap confidence intervals.